### v1

# A Differential Perspective on Distributional Reinforcement Learning

Abstract
To date, distributional reinforcement learning (distributional RL) methods have exclusively focused on the discounted setting, where an agent aims to optimize a potentially-discounted sum of rewards over time. In this work, we extend distributional RL to the average-reward setting, where an agent aims to optimize the reward received per time-step. In particular, we utilize a quantile-based approach to develop the first set of algorithms that can successfully learn and/or optimize the long-run per-step reward distribution, as well as the differential return distribution of an average-reward MDP. We derive proven-convergent tabular algorithms for both prediction and control, as well as a broader family of algorithms that have appealing scaling properties. Empirically, we find that these algorithms consistently yield competitive performance when compared to their non-distributional equivalents, while also capturing rich information about the long-run reward and return distributions.
1 Introduction
Distributional reinforcement learning (distributional RL) Morimura et al. [2010], Bellemare et al. [2017, 2023] equips decision-making agents with the ability to learn and reason about the probability distribution over a given objective. This approach transcends the traditional RL paradigm of focusing solely on expected values, thereby offering a deeper and more insightful understanding of the variability, uncertainty, and risk associated with a given objective. To date, distributional RL methods have exclusively focused on the discounted setting, where an RL agent aims to optimize a potentially-discounted sum of rewards over time (e.g. Bellemare et al. [2017], Dabney et al. [2018b, a]). In this work, we extend distributional RL to the average-reward setting, where an RL agent aims to optimize the reward received per time-step.
Importantly, this extension of distributional RL into the average-reward setting requires that we change the underlying Markov decision process (MDP) Puterman [1994] through which we operate. This means that we must move away from the discounted MDP, which forms the basis of existing distributional RL methods, and, in turn, utilize the less-explored average-reward MDP. As we will discuss in this work, changing the underlying MDP has significant implications as it relates to formulating the distributional RL approach. Most notably, we cannot rely on the widely-adopted (discounted) distributional Bellman operator Bellemare et al. [2017, 2023] in the average-reward setting. This difficulty arises because the average-reward MDP is built upon a different theoretical paradigm than the discounted MDP, thereby making it incompatible with the (discounted) distributional Bellman operator. In fact, extending distributional RL into the average-reward setting requires that we rethink and answer the following foundational questions: what do we want to learn? and how can we learn it? In this work, we address these two questions, and in the process, derive, to the best of our knowledge, the first set of distributional RL algorithms for the average-reward setting.
Our work is organized as follows: In Section 2, we provide a brief overview of related work. In Section 3, we give an overview of the fundamental concepts related to average-reward RL, quantile regression, and distributional RL. In Section 4, we motivate and derive our differential distributional RL algorithms. In Section 5, we provide empirical results which show that our algorithms consistently yield competitive performance when compared to their non-distributional equivalents, while also capturing rich information about the long-run reward and return distributions. Finally, in Section 6, we emphasize the appeal of our algorithms, highlight some of their limitations, and suggest some directions for future research.
2 Related Work
The distributional aspect of reinforcement learning has been long-studied. Early works on the subject include Sobel [1982], White [1988], and Morimura et al. [2010]. Modern distributional RL is typically associated with the distributional Bellman operator, which was introduced in Bellemare et al. [2017], along with results which showed that such methods could outperform their non-distributional equivalents in various learning tasks. Subsequent research has expanded upon this framework, incorporating novel methods that typically fall into two categories: categorical approaches Rowland et al. [2018] and quantile-based approaches Rowland et al. [2024]. While these methods have proven to be effective in the discounted setting, no attention has been given to the average-reward setting, which operates under a different theoretical framework. In this work, we address this gap by introducing the first distributional RL algorithms specifically designed for the average-reward setting.
3 Preliminaries
3.1 Average-Reward Reinforcement Learning
A finite average-reward MDP is the tuple $\mathcal{M}\doteq\langle\mathcal{S},\mathcal{A},\mathcal{R},p\rangle$, where $\mathcal{S}$ is a finite set of states, $\mathcal{A}$ is a finite set of actions, $\mathcal{R}\subset\mathbb{R}$ is a finite set of rewards, and $p:\mathcal{S}\,\times\,\mathcal{A}\,\times\,\mathcal{R}\,\times\,\mathcal{S}\rightarrow{}[0,1]$ is a probabilistic transition function that describes the dynamics of the environment. At each discrete time step, $t=0,1,2,\ldots$, an agent chooses an action, $A_{t}\in\mathcal{A}$, based on its current state, $S_{t}\in\mathcal{S}$, and receives a reward, $R_{t+1}\in\mathcal{R}$, while transitioning to a (potentially) new state, $S_{t+1}$, such that $p(s^{\prime},r\mid s,a)=\mathbb{P}(S_{t+1}=s^{\prime},R_{t+1}=r\mid S_{t}=s,A_{t}=a)$. In an average-reward MDP, an agent aims to find a policy, $\pi:\mathcal{S}\rightarrow{}\mathcal{A}$, that optimizes the long-run (or limiting) average-reward, $\bar{r}$, which is defined as follows for a given policy, $\pi$:
$\bar{r}_{\pi}(s)\doteq\lim_{n\rightarrow{}\infty}\frac{1}{n}\sum_{t=1}^{n}\mathbb{E}[R_{t}\mid S_{0}=s,A_{0:t-1}\sim\pi].$ (1)
In this work, we limit our discussion to stationary Markov policies, which are time-independent policies that satisfy the Markov property.
When working with average-reward MDPs, it is common to simplify Equation (1) into a more workable form by making certain assumptions about the Markov chain induced by following policy $\pi$. To this end, a unichain assumption is typically used when doing prediction (learning) because it ensures the existence of a unique limiting distribution of states, $\mu_{\pi}(s)\doteq\lim_{t\rightarrow{}\infty}\mathbb{P}(S_{t}=s\mid A_{0:t-1}\sim\pi)$, that is independent of the initial state, thereby simplifying Equation (1) to the following:
$\bar{r}_{\pi}=\sum_{s\in\mathcal{S}}\mu_{\pi}(s)\sum_{a\in\mathcal{A}}\pi(a\mid s)\sum_{s^{\prime}\in\mathcal{S}}\sum_{r\in\mathcal{R}}p(s^{\prime},r\mid s,a)r.$ (2)
Similarly, a communicating assumption is typically used when doing control (optimization) because it ensures the existence of a unique optimal average-reward, $\bar{r}*$, that is independent of the initial state.
The return of an MDP, $G_{t}$, captures how rewards are aggregated over the time horizon. In an average-reward MDP, the return is referred to as the differential return, and is defined as follows:
$G_{t}\doteq R_{t+1}-\bar{r}_{\pi}+R_{t+2}-\bar{r}_{\pi}+R_{t+3}-\bar{r}_{\pi}+\ldots.$ (3)
To solve an average-reward MDP, solution methods such as dynamic programming or RL are typically used in conjunction with the Bellman equation (4) for the state-value function, $v_{\pi}(s)\doteq\mathbb{E}_{\pi}[G_{t}\mid S_{t}=s]=\mathbb{E}_{\pi}[R_{t+1}-\bar{r}_{\pi}+G_{t+1}\mid S_{t}=s]$, or the Bellman optimality equation (5) for the state-action value function, $q_{\pi}(s,a)\doteq\mathbb{E}_{\pi}[G_{t}\mid S_{t}=s,A_{t}=a]$.
$v_{\pi}(s)=\sum_{a}\pi(a|s)\sum_{s^{\prime},r}p(s^{\prime},r\mid s,a)[r-\bar{r}_{\pi}+v_{\pi}(s^{\prime})]$ (4)
$q_{\pi}(s,a)=\sum_{s^{\prime},r}p(s^{\prime},r\mid s,a)[r-\bar{r}_{\pi}+\max_{a^{\prime}}q_{\pi}(s^{\prime},a^{\prime})]$ (5)
Solution methods for average-reward MDPs are typically referred to as differential methods because of the reward difference (i.e., $r-\bar{r}_{\pi}$) operation that occurs in the Bellman equations (4) and (5). We note that solution methods typically find solutions to Equations (4) and (5) up to an additive constant. This is typically not a concern, given that the relative ordering of policies is usually what is of interest.
Typically, RL-based solutions methods will learn and/or optimize the value function and average-reward simultaneously. From an algorithmic perspective, this means that the algorithms start with initial estimates (or guesses) for the value function and average-reward, then update these estimates over time, until they have learned and/or optimized these two objectives.
3.2 Quantile Regression
Quantile regression Koenker [2005] refers to the process of estimating a predetermined quantile (or percentile) of a probability distribution from samples. More specifically, for $\tau\in(0,1)$, let $F_{w}^{-1}(\tau)$ denote the $\tau^{\text{th}}$ quantile that we are trying to estimate from probability distribution $w$ (where $F_{w}:\mathbb{R}\to[0,1]$ denotes the CDF of $w$). Quantile regression maintains an estimate, $\theta$, of this value, and updates the estimate based on samples drawn from $w$ (i.e., $y\sim w$) as follows Bellemare et al. [2023]:
$\theta\xleftarrow{}\theta+\alpha_{t}(\tau-\mathds{1}_{\{y<\theta\}}),$ (6)
where $\alpha_{t}$ denotes the step size for the update. The estimate for $\theta$ will continue to adjust until the equilibrium point, $\theta^{*}$, which corresponds to $F_{w}^{-1}(\tau)$, is reached [Bellemare et al., 2023]. In other words, we have that
00 0 $=\mathbb{E}[(\tau-\mathds{1}_{\{y<\theta^{*}\}})]$ (7a)
$=\tau-\mathbb{E}[\mathds{1}_{\{y<\theta^{*}\}})]$ (7b)
$=\tau-\mathbb{P}(y<\theta^{*})$ (7c)
$\implies\theta^{*}$ $=F_{w}^{-1}(\tau).$ (7d)
3.3 Discounted Distributional Reinforcement Learning
Like average-reward MDPs, discounted MDPs have their own return which captures how rewards are aggregated over the time horizon. More specifically, let
$G^{\gamma}_{t}=\sum_{t\geq 0}{\gamma^{k}R_{t+k+1}}=R_{t+1}+\gamma G^{\gamma}_{t+1}$ (8)
denote the discounted return from a discounted MDP, where $R_{t}\in\mathcal{R}$ denotes the per-step reward and $\gamma\in[0,1]$ denotes the discount factor Puterman [1994], Sutton and Barto [2018]. The aim of discounted distributional RL is to learn the probability distribution over discounted returns. More formally, discounted distributional RL aims to learn the discounted return distribution function, $\Phi^{\gamma}_{\pi}$, such that, with a slight abuse of notation, $\Phi^{\gamma}_{\pi}(s)$ denotes the probability distribution over discounted returns when starting from state $s\in\mathcal{S}$ and following policy $\pi$, and $\Phi^{\gamma}_{\pi}(s,a)$ denotes the probability distribution over discounted returns when starting from state $s\in\mathcal{S}$, taking action $a\in\mathcal{A}$, and following policy $\pi$.
Broadly speaking, discounted distributional RL methods can be categorized based on how they approximate (or parameterize) the discounted return distribution function. In this work, we take inspiration from quantile-based methods Dabney et al. [2018b], Rowland et al. [2024], which parameterize the discounted return distribution function as follows:
$\Phi^{\gamma}_{\pi}(s)=\sum_{i=1}^{n}{\frac{1}{n}\delta_{\Omega^{\gamma}_{i}(s)}}\quad\text{or}\quad\Phi^{\gamma}_{\pi}(s,a)=\sum_{i=1}^{n}{\frac{1}{n}\delta_{\Omega^{\gamma}_{i}(s,a)}},$ (9)
where $\Omega^{\gamma}_{i}$ denotes the $\tau_{i}$-quantile of the discounted return distribution, and $\delta_{\Omega^{\gamma}_{i}}$ denotes a Dirac at $\Omega^{\gamma}_{i}$. More formally, the set of $\tau$-quantiles of a probability distribution can be defined as follows:
Definition 3.1 (Rowland et al. [2024]).
Let $\mathscr{P}(\mathbb{R})$ denote the set of probability distributions over $\mathbb{R}$. For a probability distribution $w\in\mathscr{P}(\mathbb{R})$ and parameter $\tau\in(0,1)$, the set of $\tau$-quantiles of $w$ is given by the set $\{z\in\mathbb{R}:F_{w}(z)=\tau\}\cup\text{inf}\{y\in\mathbb{R}:F_{w}(y)>\tau\}$, where $F_{w}:\mathbb{R}\to[0,1]$ is the CDF of $w$, defined by $F_{w}(t)=\mathbb{P}_{Z\sim w}(Z\leq t)$ for all $t\in\mathbb{R}$.
Quantile-based discounted distributional RL methods learn the $\tau$-quantiles of the discounted return distribution, $\{\Omega^{\gamma}_{i}\}_{i=1}^{n}$, from samples using quantile regression (see Equation 6) as follows [Dabney et al., 2018b]:
$\Omega^{\gamma}_{i,t+1}=\Omega^{\gamma}_{i,t}+\alpha_{t}\frac{1}{n}\sum_{j=1}^{n}\left[\tau_{i}-\mathds{1}_{\{\psi_{t}<0\}}\right],\;\forall i=1,2,\ldots,n,$ (10)
where $\Omega^{\gamma}_{i,t}$ denotes the estimate of the $\tau_{i}$-quantile of the discounted return distribution at time $t$, $\alpha_{t}$ denotes the step size for the update, and $\{\tau_{i}\in(0,1)\}_{i=1}^{n}$, such that $\tau_{i}=\frac{2i-1}{2n},\;i=1,2,\ldots,n$. Here, the choice of $\psi_{t}$ depends on whether we want to do prediction (learning) or control (optimization). In the case of prediction, $\psi_{t}=R_{t}+\gamma\Omega^{\gamma}_{j,t}(S_{t+1})-\Omega^{\gamma}_{i,t}(S_{t})$. In the case of control (via Q-learning Watkins and Dayan [1992]), $\psi_{t}=R_{t}+\gamma\Omega^{\gamma}_{j,t}(S_{t+1},a^{*})-\Omega^{\gamma}_{i,t}(S_{t},A_{t})$, where $a^{*}\doteq\text{argmax}_{a^{\prime}}\sum_{k=1}^{n}{\frac{1}{n}\Omega^{\gamma}_{k,t}(S_{t+1},a^{\prime})}$.
Let us take a moment to highlight the definition of the greedy action, $a^{*}$. We are being greedy with respect to the average of the $\tau$-quantiles of the discounted return distribution, which, by definition, is equivalent to being greedy with respect to the expected return, or state-action value function, $Q(s,a)$ (e.g. see Equation (5)). In a similar fashion, we can define a greedy action selection rule for our quantile-based discounted distributional RL algorithm as follows:
Definition 3.2 (Bellemare et al. [2023]).
Let $\boldsymbol{\pi}_{{}_{\text{MS}}}$ denote the space of stationary Markov policies. A greedy action selection rule is a mapping, $\mathcal{G}:\mathbb{R}^{\mathcal{S}\times\mathcal{A}}\to\boldsymbol{\pi}_{{}_{\text{MS}}}$, with the property that for any $Q\in\mathbb{R}^{\mathcal{S}\times\mathcal{A}}$, $\mathcal{G}(Q)$ is greedy with respect to $Q$. That is, $\mathcal{G}(Q)(a\mid s)>0\implies Q(s,a)=\max_{a^{\prime}\in\mathcal{A}}Q(s,a^{\prime})$.
Importantly, we note that although discounted distributional RL methods make it possible to capture information related to the underlying (discounted) return distribution, much of the theoretical guarantees associated with these methods in the control setting require an action selection rule that satisfies Definition 3.2. That is, they require a greedy action selection rule that is greedy with respect to $Q(s,a)$ [Bellemare et al., 2023].
We end this section by noting that the convergence of discounted distributional RL approaches, including the above quantile-based approach, is governed by the (discounted) distributional Bellman operator Bellemare et al. [2017, 2023]. Importantly, under certain conditions, the distributional Bellman operator is a contraction in the Wasserstein metric, thereby ensuring convergence to a fixed point. We refer the reader to Bellemare et al. [2017, 2023] for more details on the distributional Bellman operator.
4 Differential Distributional RL
In this section, we present our primary contribution: the first set of distributional RL algorithms specifically designed for the average-reward setting. We call these algorithms, Differential Distributional RL algorithms, due to the differential nature of average-reward RL algorithms.
Importantly, developing distributional RL algorithms for the average-reward setting requires that we move away from using the discounted MDP, which forms the basis of existing distributional RL methods, and, in turn, utilize the average-reward MDP. This has significant implications as it relates to formulating the distributional RL approach. Most notably, we cannot rely on the (discounted) distributional Bellman operator [Bellemare et al., 2017, 2023] because the average-reward MDP is built upon a different theoretical paradigm than the discounted MDP, thereby making it incompatible with the (discounted) distributional Bellman operator.
In this regard, we note that this situation is analogous to the non-distributional case, where non-distributional average-reward methods do not (and cannot) rely on the (regular, non-distributional) Bellman operator Puterman [1994] used in non-distributional discounted methods. Hence, like the non-distributional case, we must rely on the theory that the average-reward MDP is built upon when formulating our approach. We refer readers to texts such as Puterman [1994], Bertsekas [1995a, b], and Tijms [2003] for detailed discussions on the theoretical foundations of discounted and average-reward MDPs.
All in all, extending distributional RL into the average-reward setting requires that we rethink and answer the following foundational questions: what do we want to learn? and how can we learn it? In this section, we address these two questions, and in the process, derive the first set of distributional RL algorithms for the average-reward setting.
4.1 The Differential Distributional Objective
Before we can derive our algorithms, we first need to establish what it is that we want to learn and/or optimize. In particular, the average-reward formulation suggests two distributions that may be of interest. The first corresponds to the probability distribution over differential returns (where the differential return is defined in Equation (3)). The second corresponds to the limiting (or long-run) per-step reward distribution.
Let us first consider the probability distribution over differential returns (or the return distribution) as a distributional objective. At a first glance, this may seem like the logical distributional objective to pursue in the average-reward setting, given that its discounted analogue (i.e., the probability distribution over discounted returns) is the widely-accepted distributional objective in the discounted setting. However, this line of reasoning ignores a fundamental difference between the discounted and average-reward settings.
In particular, first consider the discounted setting: taking the mean of the probability distribution over discounted returns yields the expected (discounted) return, which is the non-distributional objective in the discounted setting. As such, it is easy to argue that the probability distribution over discounted returns is the natural choice for the distributional objective in the discounted setting, given that the discounted return is both the explicit and implicit learning and/or optimization objective in this setting.
Now consider the average-reward setting: taking the mean of the probability distribution over differential returns yields the expected (differential) return, which is not the non-distributional objective in the average-reward setting; the long-run average-reward is (i.e., Equation (1)). In other words, although the differential return (see Equation (3)) serves as the explicit learning and/or optimization objective in the average-reward setting, the implicit and ultimately primary objective is the long-run average-reward (as defined in Equation (1)). As such, it is more difficult to argue that the probability distribution over differential returns is the natural choice for the distributional objective in the average-reward setting.
Let us now consider the limiting (or long-run) per-step reward distribution as a distributional objective for the average-reward setting. Taking the mean of the limiting per-step reward distribution yields the long-run average-reward, which indeed corresponds to the non-distributional objective in the average-reward setting. As such, the limiting per-step reward distribution appears to be a much more appealing and natural distributional objective for the average-reward setting.
Hence, given the above reasoning, which is formalized as Proposition 4.1 below, the primary focus of this work will be to derive differential algorithms that can learn and/or optimize the limiting per-step reward distribution. We will briefly revisit the return distribution from a purely empirical perspective in Section 4.3.
Proposition 4.1.
The limiting per-step reward distribution is the natural distributional objective in the average-reward setting, given that its mean yields the long-run average-reward, which is the primary prediction and control objective of average-reward reinforcement learning.
We are now ready to begin our theoretical treatment of the limiting per-step reward distribution. We begin by formally defining our distributional objective. In particular, for a given policy, $\pi$, let
$\phi_{\pi}(s)\doteq\lim_{t\rightarrow{}\infty}\mathbb{P}(R_{t}\mid S_{0}=s,A_{0:t-1}\sim\pi)$ (11)
denote the limiting per-step reward distribution induced by following policy $\pi$, where $R_{t}\in\mathcal{R}$ denotes the per-step reward. As is standard practice with the regular average-reward objective, we can simplify Equation (11) by making certain assumptions about the Markov chain induced by following policy $\pi$. To this end, we will utilize a unichain assumption when doing prediction (learning), because it ensures the existence of a unique limiting distribution of states that is independent of the initial state, such that $\phi_{\pi}(s)=\phi_{\pi}$. Similarly, we will utilize a communicating assumption when doing control (optimization), because it ensures the existence of a unique optimal average-reward, along with a corresponding per-step reward distribution that is independent of the initial state. Both assumptions are formally listed below:
Assumption 4.2 (Unichain Assumption for Prediction).
The Markov chain induced by the policy is unichain. That is, the induced Markov chain consists of a single recurrent class and a potentially-empty set of transient states.
Assumption 4.3 (Communicating Assumption for Control).
The induced Markov chain has a single communicating class. That is, each state in the induced Markov chain is accessible from every other state under a deterministic stationary policy.
Figure 1 depicts the agent-environment interaction in an average-reward MDP, where following policy $\pi$ yields a limiting reward distribution, $\phi_{\pi}$, whose mean corresponds to the average-reward, $\bar{r}_{\pi}$.
Figure 1: Illustration of the agent-environment interaction in an average-reward MDP. As $t\to\infty$, following policy $\pi$ yields a limiting per-step reward distribution, $\phi_{\pi}$, with an average-reward, $\bar{r}_{\pi}$. Standard average-reward RL methods aim to learn and/or optimize the average-reward, $\bar{r}_{\pi}$. By contrast, differential distributional RL aims to learn and/or optimize the limiting per-step reward distribution, $\phi_{\pi}$.
4.2 A Quantile-Based Approach
In this section, we take inspiration from the quantile-based methods utilized for discounted distributional RL to develop a quantile-based approach that can be used to learn and/or optimize the limiting per-step reward distribution. Like the discounted setting (see Equation (10)), our overall goal and strategy is to derive an appropriate set of quantile regression-based updates that approximate our target distribution.
More formally, we approximate (or parameterize) the limiting per-step reward distribution as follows:
$\phi_{\pi}=\sum_{i=1}^{m}{\frac{1}{m}\delta_{\theta_{i}}},$ (12)
where $\theta_{i}$ denotes the $\tau_{i}$-quantile of the limiting per-step reward distribution, and $\delta_{\theta_{i}}$ denotes a Dirac at $\theta_{i}$. Here, we adopt the same formal definition for the $\tau$-quantiles of a probability distribution as in the discounted setting (i.e., Definition 3.1).
Now, let us consider the $\tau$-quantiles of the limiting per-step reward distribution, $\{\theta_{i}\}_{i=1}^{m}$. As previously mentioned, we wish to approximate these quantiles using quantile regression. Looking at the generic quantile regression update rule (6), we can derive an appropriate set of quantile update rules for the per-step reward quantiles, as follows:
$\theta_{i,t+1}=\theta_{i,t}+\alpha_{t}\left(\tau_{i}-\mathds{1}_{\{R_{t}<\theta_{i,t}\}}\right)\;\forall i=1,2,\ldots,m,$ (13)
where $\theta_{i,t}$ denotes the estimate of the $\tau_{i}$-quantile of the limiting per-step reward distribution, $\alpha_{t}$ denotes the step size for the update, and $R_{t}$ denotes the per-step reward.
When comparing this quantile update rule to that of the discounted setting (i.e., Equation (10)), we can see that both rules have a similar structure, with the key difference being that we have replaced $\psi_{t}$ with $R_{t}-\theta_{i,t}$. Importantly, the per-step reward quantile update rule (13) does not contain the full TD or Q-learning targets and estimates. This is not necessarily an issue, so long as we can find a way to properly incorporate the quantile update rule (13) into a broader algorithm that does contain the full TD or Q-learning targets and estimates.
To this end, let us consider a typical average-reward RL algorithm in the non-distributional setting. As previously mentioned, such an algorithm typically learns and/or optimizes the value function and average-reward simultaneously. From an algorithmic perspective, this means that we start with initial guesses for the value function and average-reward, then update these estimates over time, until we have learned and/or optimized these two objectives. For instance, a subset of average-reward RL algorithms aims to learn both objectives using the temporal-difference (or TD) error (e.g. Schwartz [1993], Wan et al. [2021], Rojas and Lee [2025]). Below is an example (tabular) TD-learning algorithm:
$\delta_{t}=R_{t+1}-\bar{R}_{t}+V_{t}(S_{t+1})-V_{t}(S_{t})$ (14a)
$V_{t+1}(S_{t})=V_{t}(S_{t})+\alpha_{t}\delta_{t}$ (14b)
$\bar{R}_{t+1}=\bar{R}_{t}+\eta\alpha_{t}\delta_{t}$ (14c)
where, $V_{t}:\mathcal{S}\rightarrow\mathbb{R}$ is a table of state-value function estimates, $\alpha_{t}$ is the step size, $\delta_{t}$ is the TD error, $\bar{R}_{t}$ is an estimate of the average-reward, $\bar{r}_{\pi}$, and $\eta$ is a positive scalar.
Hence, our goal is to somehow incorporate the quantile updates (13) into an overall differential algorithm, like the one shown in Equation (14). In this work, we explore two such approaches that achieve this. We discuss the first of these two approaches below. We discuss the second approach in Section 4.3.
We now discuss the first approach, through which we incorporate the quantile updates into a differential algorithm by replacing the average-reward update (14c) with a set of equivalent, quantile-based updates. In particular, we can relate the average-reward, $\bar{r}_{\pi}$, to the $\tau$-quantiles of the limiting per-step reward distribution, $\{\theta_{i}\}_{i=1}^{m}$, as follows:
Lemma 4.4.
Given assumptions 4.2 and 4.3, consider the limiting per-step reward distribution parameterized by the $\tau$-quantiles, $\{\theta_{i}\}_{i=1}^{m}$, as per Equation (12) for $\{\tau_{i}\in(0,1)\}_{i=1}^{m}$, such that $\tau_{i}=\frac{2i-1}{2m},\;i=1,2,\ldots,m$. The expected value of the $\tau$-quantiles converges to the average-reward of the limiting per-step reward distribution, $\bar{r}_{\pi}$, as $m\to\infty$.
Proof.
Let $F(r)$ denote the CDF of the limiting per-step reward distribution (which exists and is unique given assumptions 4.2 and 4.3), and let $\theta_{i}$ represent its $\tau_{i}$-quantile, such that $F(\theta_{i})=\tau_{i}$. The expected value of the $\tau$-quantiles is:
$\frac{1}{m}\sum_{i=1}^{m}\theta_{i}=\frac{1}{m}\sum_{i=1}^{m}F^{-1}\left(\frac{2i-1}{2m}\right).$ (15)
By the definition of $\tau_{i}=\frac{2i-1}{2m}$, the $\tau_{i}$ values are evenly spaced over $[0,1]$, creating a uniform partition. Hence, as $m\to\infty$, the summation $\frac{1}{m}\sum_{i=1}^{m}\theta_{i}$ becomes a Riemann sum for the following integral:
[…]
Thus, the expected value of the $\tau$-quantiles converges to the average reward, $\bar{r}_{\pi}$, of the limiting per-step reward distribution as $m\to\infty$. This completes the proof. ∎
Hence, we now have a way to express the set of quantile updates (13) as an average-reward update. In particular, we can replace the average-reward update (14c) with a set of equivalent, quantile-based updates as follows:
$\theta_{i,t+1}$ $=\theta_{i,t}+\alpha_{i,t}\left(\tau_{i}-\mathds{1}_{\{R_{t+1}<\theta_{i,t}\}}\right),\;\forall i=1,2,\ldots,m,$ (18a)
$\bar{R}_{t+1}$ $=\frac{1}{m}\sum_{i=1}^{m}\theta_{i,t+1}.$ (18b)
As such, by replacing the average-reward update (14c) with the equivalent set of quantile-based updates (18), we now have a way to learn the limiting per-step reward distribution. Importantly, we can show that this equivalent set of quantile-based updates converges to the average-reward of the limiting per-step reward distribution induced by a given policy, $\pi$:
Theorem 4.5.
Given assumptions 4.2 and 4.3, consider the limiting per-step reward distribution induced by following policy $\pi$, and parameterized by the $\tau$-quantiles, $\{\theta_{i}\}_{i=1}^{m}$, as per Equation (12) for $\{\tau_{i}\in(0,1)\}_{i=1}^{m}$, such that $\tau_{i}=\frac{2i-1}{2m},\;i=1,2,\ldots,m$. Also consider the quantile-based update rules for the average-reward estimate (18), along with corresponding step sizes that satisfy: $\alpha_{t}>0,\;\sup_{t\geq 0}\alpha_{t}<\infty,\;\sum_{t=0}^{\infty}\alpha_{t}=\infty,\text{and}\;\alpha_{t}=o(1/\text{log }t)$. If the quantile estimates, $\{\theta_{i,t}\}_{i=1}^{m}$, converge, almost surely, to the $\tau$-quantiles of the limiting per-step reward distribution, then the average-reward estimate, $\bar{R}_{t}$, converges, almost surely, to the average-reward of the limiting per-step reward distribution, $\bar{r}_{\pi}$, as $t\to\infty$.
Proof.
If $\{\theta_{i,t}\}_{i=1}^{m}\to\{\theta_{i}\}_{i=1}^{m}$ as $t\to\infty$, then, by Lemma 4.4, it follows directly that the average of these estimates, which corresponds to $\bar{R}_{t}$, converges, almost surely, to $\bar{r}_{\pi}$ as $t\to\infty$. This completes the proof. ∎
Hence, given the convergence of the $\tau$-quantile estimates, it is intuitive to see that the average-reward estimate will also converge. We provide a full convergence proof for the $\tau$-quantile estimates in Appendix B. At a high level, the aforementioned proof shows that because the quantile-based updates (18a) are of the same form as Equation (6), the estimates will converge to the desired $\tau$-quantiles via the logic outlined by Equation (7).
As such, we now have a way to learn the limiting per-step reward distribution. We can extend these results into the control setting by choosing an appropriate action selection rule. As in the discounted setting, we will utilize an action selection rule that is consistent with Definition 3.2. To this end, Algorithm 1 shows the Differential Distributional (or D2) Q-learning algorithm, which adopts a greedy action selection rule with respect to the state-action value function. The full set of Differential Distributional RL algorithms is included in Appendix A.
Algorithm 1 Differential Distributional (D2) Q-Learning (Tabular)
  Obtain initial $S$
  while still time to train do
     $A\leftarrow$ action given by $\pi$ for $S$
     Take action $A$, observe $R,S^{\prime}$
     $\bar{R}=\frac{1}{m}\sum_{i=1}^{m}\theta_{i}$
     $\delta=R-\bar{R}+\max_{a}Q(S^{\prime},a)-Q(S,A)$
     $Q(S,A)=Q(S,A)+\alpha\delta$
     for $i=1,2,\ldots,m$ do
        $\theta_{i}=\theta_{i}+\alpha_{\theta}\left(\tau_{i}-\mathds{1}_{\{R<\theta_{i}\}}\right)$
     $S=S^{\prime}$
  return $\{\theta_{i}\}_{i=1}^{m}$
Importantly, with the aforementioned action selection rule, we retain the regular (non-distributional) Q-learning operator for the average-reward setting:
$TQ(s,a)\doteq\sum_{s^{\prime},r}p(s^{\prime},r\;|\;s,a)(r+\max_{a^{\prime}}Q(s^{\prime},a^{\prime})).$ (19)
As such, we can use existing results to establish that the convergence of the state-action value estimates implies that the average-reward converges to the optimal average-reward:
Theorem 4.6.
Let $\bar{r}_{\pi_{t}}$ denote the average-reward obtained when following policy $\pi_{t}$, such that $\pi_{t}$ is a greedy policy with respect to the state-action value estimates, $Q_{t}(s,a)$, where $s\in\mathcal{S}$ and $a\in\mathcal{A}$. Also, let $q^{*}$ denote the solution of the Bellman optimality equation (5). If assumption 4.3 holds, and there exists a unique solution of $q^{*}$ only up to a constant, then we have that if $Q_{t}\to q^{*}$ a.s. as $t\to\infty$, then $\bar{r}_{\pi_{t}}\to\bar{r}*$ a.s. as $t\to\infty$.
Proof.
This follows directly from Theorem 8.5.5 of Puterman [1994]. In particular, we have that:
$\min_{s,a}(TQ_{t}-Q_{t})$ $\leq\bar{r}_{\pi_{t}}$
$\leq\bar{r}*$
$\leq\max_{s,a}(TQ_{t}-Q_{t}),$ (20)
$\implies|\bar{r}*-\bar{r}_{\pi_{t}}|$ $\leq sp(TQ_{t}-Q_{t}),$ (21)
where $TQ(s,a)$ is defined in Equation (19).
Because $Q_{t}\to q^{*}$ a.s. as $t\to\infty$, and $sp(TQ_{t}-Q_{t})$ is clearly a continuous function of $Q_{t}$, we have, by the continuous mapping theorem, that $sp(TQ_{t}-Q_{t})\to sp(Tq^{*}-q^{*})=0$ a.s. as $t\to\infty$.
Hence, we can conclude that $\bar{r}_{\pi_{t}}\to\bar{r}_{*}$ a.s. as $t\to\infty$. ∎
The above theorem requires that $Q_{t}\to q^{*}$ a.s. as $t\to\infty$. In this regard, we provide a full convergence proof for the convergence of $Q_{t}$ in Appendix B. We also provide a full proof for the convergence of the value function estimates in the prediction setting (i.e., $V_{t}$) in Appendix B. Altogether, the theoretical results provided in this section and in Appendix B show the almost sure convergence of the tabular D2 algorithms in both the prediction and control settings.
4.3 Learning the Differential Return Distribution
Although we argue in this work that the limiting per-step reward distribution is the natural distributional objective in the average-reward setting, it may still be useful to consider the probability distribution over (differential) returns. For instance, such an approach could potentially yield superior empirical performance, as is the case in the discounted setting (e.g. Bellemare et al. [2017], Dabney et al. [2018b, a]). As such, in this work we explore such an approach from a purely empirical perspective, and propose a set of quantile-based algorithms that parameterize the probability distribution over differential returns, as well as the limiting per-step reward distribution. We call the resulting set of algorithms, Double Differential Distributional, or D3, algorithms because they provide a means for simultaneously learning and/or optimizing both distributions. In particular, we parameterize the differential return distribution as follows:
$\Phi_{\pi}(s)=\sum_{i=1}^{n}{\frac{1}{n}\delta_{\Omega_{i}(s)}}\quad\text{or}\quad\Phi_{\pi}(s,a)=\sum_{i=1}^{n}{\frac{1}{n}\delta_{\Omega_{i}(s,a)}},$ (22)
where $\Phi_{\pi}$ denotes the probability distribution over differential returns (when starting from state $s\in\mathcal{S}$ or when starting from state $s\in\mathcal{S}$ and taking action $a\in\mathcal{A}$, and following policy $\pi$), $\Omega_{i}$ denotes the $\tau_{i}$-quantile of the differential return distribution, and $\delta_{\Omega_{i}}$ denotes a Dirac at $\Omega_{i}$. Our aim is to learn the $\tau$-quantiles of the differential return distribution, $\{\Omega_{i}\}_{i=1}^{n}$, from samples using quantile regression as follows:
$\Omega_{i,t+1}=\Omega_{i,t}+\alpha_{t}\frac{1}{n}\sum_{j=1}^{n}\left[\tau_{i}-\mathds{1}_{\{\psi_{t}<0\}}\right],\;\forall i=1,2,\ldots,n,$ (23)
where $\Omega_{i,t}$ denotes the estimate of the $\tau_{i}$-quantile of the differential return distribution at time $t$, $\alpha_{t}$ denotes the step size for the update, and $\{\tau_{i}\in(0,1)\}_{i=1}^{n}$, such that $\tau_{i}=\frac{2i-1}{2n},\;i=1,2,\ldots,n$. Here, the choice of $\psi_{t}$ depends on whether we want to do prediction (learning) or control (optimization). In the case of prediction, $\psi_{t}=R_{t}-\bar{R}_{t}+\Omega_{j,t}(S_{t+1})-\Omega_{i,t}(S_{t})$, where $\bar{R}_{t}$ is the average-reward estimate. In the case of control (via Q-learning), $\psi_{t}=R_{t}-\bar{R}_{t}+\Omega_{j,t}(S_{t+1},a^{*})-\Omega_{i,t}(S_{t},A_{t})$, where $a^{*}\doteq\text{argmax}_{a^{\prime}}\sum_{k=1}^{n}{\frac{1}{n}\Omega_{k,t}(S_{t+1},a^{\prime})}$.
Algorithm 2 shows the Double Differential Distributional (or D3) Q-learning algorithm, which adopts a greedy action selection rule with respect to the average of the $\tau$-quantiles of the differential return distribution. The full set of Double Differential Distributional RL algorithms is included in Appendix A.
Algorithm 2
Double Differential Distributional (D3) Q-Learning (Tabular)
  Obtain initial $S$
  while still time to train do
     $A\leftarrow$ action given by $\pi$ for $S$
     Take action $A$, observe $R,S^{\prime}$
     $\bar{R}=\frac{1}{m}\sum_{i=1}^{m}\theta_{i}$
     $a^{*}=\text{argmax}_{a^{\prime}}\sum_{j=1}^{n}{\frac{1}{n}\Omega_{j}(S^{\prime},a^{\prime})}$
     for $j=1,2,\ldots,n$ do
        $\Omega_{j}(S,A)=\Omega_{j}(S,A)+\alpha\frac{1}{n}\sum_{k=1}^{n}\Big{[}\tau_{j}-\mathds{1}_{\{R-\bar{R}+\Omega_{k}(S^{\prime},a^{*})-\Omega_{j}(S,A)<0\}}\Big{]}$
     for $i=1,2,\ldots,m$ do
        $\theta_{i}=\theta_{i}+\alpha_{\theta}\left(\tau_{i}-\mathds{1}_{\{R<\theta_{i}\}}\right)$
     $S=S^{\prime}$
  return $\{\theta_{i}\}_{i=1}^{m}$ and $\{\Omega_{j}\}_{j=1}^{n}$
5 Experimental Results
In this section, we present empirical results obtained when applying our D2 and D3 algorithms on two RL tasks. These two tasks were chosen because, in both cases, the optimal limiting per-step reward distribution is known, thereby making it possible to directly gauge the accuracy of the learned distributions from our algorithms. The first task is the two-state red-pill blue-pill environment [Rojas and Lee, 2025], and the second task is the well-known inverted pendulum environment. We compared the performance of our algorithms to that of non-distributional differential algorithms derived from the framework proposed in Wan et al. [2021].
In terms of empirical results, Figures 2 and 3 show the agent’s (per-step) reward quantile estimates as learning progresses when using the D2 Q-learning algorithm in the red-pill blue-pill and inverted pendulum tasks, respectively. As shown in the figures, in both tasks, the agent’s quantile estimates converge to the quantiles of the limiting per-step reward distribution induced by the optimal policy, thereby empirically showing that the D2 algorithm not only converges, but converges to the optimal solution. In addition, these results highlight an important property of our quantile-based approach: the average-reward estimate converges to the actual average-reward induced by a given policy. This is in contrast to many prior average-reward RL algorithms, whose average-reward estimates converge to the actual average-reward up to an additive constant.
Figure 2: a) Histogram showing the empirical ($\varepsilon$-greedy) optimal (long-run) per-step reward distribution in the red-pill blue-pill task. b) Quantiles of the ($\varepsilon$-greedy) optimal (long-run) per-step reward distribution in the red-pill blue-pill task. c) Convergence plot of the agent’s (per-step) reward quantile estimates as learning progresses when using the D2 Q-learning algorithm in the red-pill blue-pill task.
Figure 3: a) Histogram showing the empirical ($\varepsilon$-greedy) optimal (long-run) per-step reward distribution in the inverted pendulum task. b) Quantiles of the ($\varepsilon$-greedy) optimal (long-run) per-step reward distribution in the inverted pendulum task. c) Convergence plot of the agent’s (per-step) reward quantile estimates as learning progresses when using the D2 actor-critic algorithm in the inverted pendulum task.
Figure 4 shows the rolling average-reward as learning progresses when using the D2 and D3 algorithms vs. the non-distributional Differential algorithm in the two tasks. As shown in the figure, in both tasks, the D2 and D3 algorithms consistently yield competitive performance when compared to their non-distributional counterpart. We note that the high-dimensionality of the state-space for the inverted pendulum environment required that we use function approximation for this task. Hence, the results in Figure 4 also show that our quantile-based approach yields competitive performance in both tabular and function approximation-based settings. The full set of experimental details and results can be found in Appendix C.
Figure 4: Rolling average-reward when using the D2 and D3 algorithms vs. the (non-distributional) Differential algorithm in the a) red-pill blue-pill, and b) inverted pendulum environments. A solid line denotes the mean average-reward, and the corresponding shaded region denotes a 95% confidence interval.
6 Discussion, Limitations, and Future Work
In this work, we introduced the first distributional RL algorithms specifically designed for the average-reward setting. In particular, we defined an appropriate distributional objective for the average-reward setting, as well as derived two quantile-based approaches that can learn and/or optimize this new distributional objective. We showed, both theoretically and empirically, that these Differential Distributional algorithms are able to learn the average-reward-optimal policy, as well as the corresponding (optimal) distributional objective.
We now discuss some of the benefits and advantages enjoyed by our proposed algorithms. Firstly, it turns out that our choice of distributional objective for the average-reward setting allows the resulting D2 algorithms to be more scalable in comparison to distributional RL algorithms in the discounted setting. In particular, the D2 algorithms only require $m$ parameters to parameterize the distributional objective, where $m$ is the number of quantiles. This is in contrast to discounted distributional algorithms, which require $|\mathcal{S}|\times|\mathcal{A}|\times m$ parameters to parameterize the discounted distributional objective.
Similarly, our algorithms enjoy reduced bias in their estimations when compared to some prior methods. In particular, some prior algorithms in the average-reward setting are only guaranteed to find the solution of the Bellman equations (4) and (5) up to an additive constant (e.g. Wan et al. [2021], Rojas and Lee [2025]). By contrast, the theoretical results provided in this work (for tabular algorithms) ensure that the average-reward estimate from our algorithm converges to the actual average-reward induced by a given policy. Consequently, the solutions to the Bellman equations (4) and (5) obtained by our algorithms correspond to the actual value functions, rather than potentially-shifted estimates.
Finally, as evidenced by the empirical results presented in this work, our algorithms are able to achieve competitive performance with non-distributional differential algorithms, while also accurately capturing rich information about the long-run reward and return distributions. Interestingly, the D3 algorithms did not outperform the D2 algorithms in either experiment. Additional empirical evaluations should be performed (with an emphasis on more complex RL tasks) to see if a consistent trend emerges.
Future work should look to address the aforementioned empirical questions, as well as tackle the theoretical questions that remain. For example, it remains to be seen how different aspects (beyond the mean) of the chosen distributional objective can be optimized, as well as how a categorical approach could be employed (rather than a quantile-based approach). Exploring these directions may yield deeper insights, and in the process, continue unlocking the full potential of distributional approaches in the average-reward setting.
Appendix A Differential Distributional RL Algorithms
In this appendix, we provide pseudocode for our differential distributional RL algorithms.
We note that, as is common practice in the discounted setting, we adopt the quantile Huber loss [Dabney et al., 2018b], $h_{\tau}^{\lambda}(x)$, in the D3 algorithms that utilize function approximation (that is, Algorithms 10, 11, and 12):
$h_{\tau}^{\lambda}(x)=\begin{cases}|\tau-\mathds{1}{\{x<0\}}|\frac{1}{2}x^{2},&\text{if }|x|\leq\lambda\\ |\tau-\mathds{1}{\{x<0\}}|\lambda\left(|x|-\frac{1}{2}\lambda\right),&\text{if}|x|>\lambda\text{.}\end{cases}$ (A.1)
Algorithm 3 Differential Distributional (D2) TD-Learning (Tabular)
  Input: the policy $\pi$ to be used (e.g., $\varepsilon$-greedy), number of quantiles, $m$, and $\tau$-locations, such that $\tau_{i}=\frac{2i-1}{2m}$
  Algorithm parameters: step size parameters $\alpha$, $\eta_{{}_{\theta}}$
  Initialize $V(s)\>\forall s$ arbitrarily (e.g. to zero)
  Initialize quantiles $\theta_{1},\theta_{2},\ldots,\theta_{m}$ arbitrarily (e.g. to zero)
  Obtain initial $S$
  while still time to train do
     $A\leftarrow$ action given by $\pi$ for $S$
     Take action $A$, observe $R,S^{\prime}$
     $\bar{R}=\frac{1}{m}\sum_{i=1}^{m}\theta_{i}$
     $\delta=R-\bar{R}+V(S^{\prime})-V(S)$
     $V(S)=V(S)+\alpha\delta$
     for $i=1,2,\ldots,m$ do
        $\theta_{i}=\theta_{i}+\eta_{{}_{\theta}}\alpha\left(\tau_{i}-\mathds{1}{\{R<\theta_{i}\}}\right)$
     $S=S^{\prime}$
  return $\{\theta_{i}\}_{i=1}^{m}$
Algorithm 4 Differential Distributional (D2) Q-Learning (Tabular)
  Input: the policy $\pi$ to be used (e.g., $\varepsilon$-greedy), number of quantiles, $m$, and $\tau$-locations, such that $\tau_{i}=\frac{2i-1}{2m}$
  Algorithm parameters: step size parameters $\alpha$, $\eta_{{}_{\theta}}$
  Initialize $Q(s,a)\>\forall s,a$ arbitrarily (e.g. to zero)
  Initialize quantiles $\theta_{1},\theta_{2},\ldots,\theta_{m}$ arbitrarily (e.g. to zero)
  Obtain initial $S$
  while still time to train do
     $A\leftarrow$ action given by $\pi$ for $S$
     Take action $A$, observe $R,S^{\prime}$
     $\bar{R}=\frac{1}{m}\sum_{i=1}^{m}\theta_{i}$
     $\delta=R-\bar{R}+\max_{a}Q(S^{\prime},a)-Q(S,A)$
     $Q(S,A)=Q(S,A)+\alpha\delta$
     for $i=1,2,\ldots,m$ do
        $\theta_{i}=\theta_{i}+\eta_{{}_{\theta}}\alpha\left(\tau_{i}-\mathds{1}{\{R<\theta_{i}\}}\right)$
     $S=S^{\prime}$
  return $\{\theta_{i}\}_{i=1}^{m}$
Algorithm 5 Differential Distributional (D2) TD-Learning (Function Approximation)
  Input: a differentiable state-value function parameterization: $\hat{v}(s,\boldsymbol{w})$, the policy $\pi$ to be used (e.g., $\varepsilon$-greedy), number of quantiles, $m$, and $\tau$-locations, such that $\tau_{i}=\frac{2i-1}{2m}$
  Algorithm parameters: step size parameters $\alpha$, $\eta_{{}_{\theta}}$
  Initialize state-value weights $\boldsymbol{w}\in\mathbb{R}^{d}$ arbitrarily (e.g. to $\boldsymbol{0}$)
  Initialize quantiles $\theta_{1},\theta_{2},\ldots,\theta_{m}$ arbitrarily (e.g. to zero)
  Obtain initial $S$
  while still time to train do
     $A\leftarrow$ action given by $\pi$ for $S$
     Take action $A$, observe $R,S^{\prime}$
     $\bar{R}=\frac{1}{m}\sum_{i=1}^{m}\theta_{i}$
     $\delta=R-\bar{R}+\hat{v}(S^{\prime},\boldsymbol{w})-\hat{v}(S,\boldsymbol{w})$
     $\boldsymbol{w}=\boldsymbol{w}+\alpha\delta\nabla\hat{v}(S,\boldsymbol{w})$
     for $i=1,2,\ldots,m$ do
        $\theta_{i}=\theta_{i}+\eta_{{}_{\theta}}\alpha\left(\tau_{i}-\mathds{1}{\{R<\theta_{i}\}}\right)$
     $S=S^{\prime}$
  return $\{\theta_{i}\}_{i=1}^{m}$
Algorithm 6 Differential Distributional (D2) Q-Learning (Function Approximation)
  Input: a differentiable state-action value function parameterization: $\hat{q}(s,a,\boldsymbol{w})$, the policy $\pi$ to be used (e.g., $\varepsilon$-greedy), number of quantiles, $m$, and $\tau$-locations, such that $\tau_{i}=\frac{2i-1}{2m}$
  Algorithm parameters: step size parameters $\alpha$, $\eta_{{}_{\theta}}$
  Initialize state-action value weights $\boldsymbol{w}\in\mathbb{R}^{d}$ arbitrarily (e.g. to $\boldsymbol{0}$)
  Initialize quantiles $\theta_{1},\theta_{2},\ldots,\theta_{m}$ arbitrarily (e.g. to zero)
  Obtain initial $S$
  while still time to train do
     $A\leftarrow$ action given by $\pi$ for $S$
     Take action $A$, observe $R,S^{\prime}$
     $\bar{R}=\frac{1}{m}\sum_{i=1}^{m}\theta_{i}$
     $\delta=R-\bar{R}+\max_{a}\hat{q}(S^{\prime},a,\boldsymbol{w})-\hat{q}(S,A,\boldsymbol{w})$
     $\boldsymbol{w}=\boldsymbol{w}+\alpha\delta\nabla\hat{q}(S,A,\boldsymbol{w})$
     for $i=1,2,\ldots,m$ do
        $\theta_{i}=\theta_{i}+\eta_{{}_{\theta}}\alpha\left(\tau_{i}-\mathds{1}{\{R<\theta_{i}\}}\right)$
     $S=S^{\prime}$
  return $\{\theta_{i}\}_{i=1}^{m}$
Algorithm 7 Differential Distributional (D2) Actor-Critic
  Input: a differentiable state-value function parameterization $\hat{v}(s,\boldsymbol{w})$, a differentiable policy parameterization $\pi(a\mid s,\boldsymbol{u})$, number of quantiles, $m$, and $\tau$-locations, such that $\tau_{i}=\frac{2i-1}{2m}$
  Algorithm parameters: step size parameters $\alpha$, $\eta_{\pi}$, $\eta_{{}_{\theta}}$
  Initialize state-value weights $\boldsymbol{w}\in\mathbb{R}^{d}$ and policy weights $\boldsymbol{u}\in\mathbb{R}^{d^{\prime}}$ (e.g. to $\boldsymbol{0}$)
  Initialize quantiles $\theta_{1},\theta_{2},\ldots,\theta_{m}$ arbitrarily (e.g. to zero)
  Obtain initial $S$
  while still time to train do
     $A\sim\pi(\cdot\mid S,\boldsymbol{u})$
     Take action $A$, observe $R,S^{\prime}$
     $\bar{R}=\frac{1}{m}\sum_{i=1}^{m}\theta_{i}$
     $\delta=R-\bar{R}+\hat{v}(S^{\prime},\boldsymbol{w})-\hat{v}(S,\boldsymbol{w})$
     $\boldsymbol{w}=\boldsymbol{w}+\alpha\delta\nabla\hat{v}(S,\boldsymbol{w})$
     $\boldsymbol{u}=\boldsymbol{u}+\eta_{\pi}\alpha\delta\nabla\text{ln}\pi(A\mid S,\boldsymbol{u})$
     for $i=1,2,\ldots,m$ do
        $\theta_{i}=\theta_{i}+\eta_{{}_{\theta}}\alpha\left(\tau_{i}-\mathds{1}{\{R<\theta_{i}\}}\right)$
     $S=S^{\prime}$
  return $\{\theta_{i}\}_{i=1}^{m}$
Algorithm 8 Double Differential Distributional (D3) TD-Learning (Tabular)
  Input: the policy $\pi$ to be used (e.g., $\varepsilon$-greedy), number of per-step reward quantiles, $m$, number of differential return quantiles, $n$, and $\tau$-locations, such that $\tau_{i}=\frac{2i-1}{2m}$ and $\tau_{j}=\frac{2j-1}{2n}$
  Algorithm parameters: step size parameters $\alpha$, $\eta_{{}_{\theta}}$
  Initialize $\Omega_{j}(s)\>\forall s,j$ arbitrarily (e.g. to zero)
  Initialize per-step reward quantiles $\theta_{1},\theta_{2},\ldots,\theta_{m}$ arbitrarily (e.g. to zero)
  Obtain initial $S$
  while still time to train do
     $A\leftarrow$ action given by $\pi$ for $S$
     Take action $A$, observe $R,S^{\prime}$
     $\bar{R}=\frac{1}{m}\sum_{i=1}^{m}\theta_{i}$
     for $j=1,2,\ldots,n$ do
        $\Omega_{j}(S)=\Omega_{j}(S)+\alpha\frac{1}{n}\sum_{k=1}^{n}\Big{[}\tau_{j}-\mathds{1}{\{R-\bar{R}+\Omega_{k}(S^{\prime})-\Omega_{j}(S)<0\}}\Big{]}$
     for $i=1,2,\ldots,m$ do
        $\theta_{i}=\theta_{i}+\eta_{{}_{\theta}}\alpha\left(\tau_{i}-\mathds{1}{\{R<\theta_{i}\}}\right)$
     $S=S^{\prime}$
  return $\{\theta_{i}\}_{i=1}^{m}$ and $\{\Omega_{j}\}_{j=1}^{n}$
Algorithm 9 Double Differential Distributional (D3) Q-Learning (Tabular)
  Input: the policy $\pi$ to be used (e.g., $\varepsilon$-greedy), number of per-step reward quantiles, $m$, number of differential return quantiles, $n$, and $\tau$-locations, such that $\tau_{i}=\frac{2i-1}{2m}$ and $\tau_{j}=\frac{2j-1}{2n}$
  Algorithm parameters: step size parameters $\alpha$, $\eta_{{}_{\theta}}$
  Initialize $\Omega_{j}(s,a)\>\forall s,a,j$ arbitrarily (e.g. to zero)
  Initialize per-step reward quantiles $\theta_{1},\theta_{2},\ldots,\theta_{m}$ arbitrarily (e.g. to zero)
  Obtain initial $S$
  while still time to train do
     $A\leftarrow$ action given by $\pi$ for $S$
     Take action $A$, observe $R,S^{\prime}$
     $\bar{R}=\frac{1}{m}\sum_{i=1}^{m}\theta_{i}$
     $a^{*}=\text{argmax}_{a^{\prime}}\sum_{j=1}^{n}{\frac{1}{n}\Omega_{j}(S^{\prime},a^{\prime})}$
     for $j=1,2,\ldots,n$ do
        $\Omega_{j}(S,A)=\Omega_{j}(S,A)+\alpha\frac{1}{n}\sum_{k=1}^{n}\Big{[}\tau_{i}-\mathds{1}{\{R-\bar{R}+\Omega_{k}(S^{\prime},a^{*})-\Omega_{j}(S,A)<0\}}\Big{]}$
     for $i=1,2,\ldots,m$ do
        $\theta_{i}=\theta_{i}+\eta_{{}_{\theta}}\alpha\left(\tau_{i}-\mathds{1}{\{R<\theta_{i}\}}\right)$
     $S=S^{\prime}$
  return $\{\theta_{i}\}_{i=1}^{m}$ and $\{\Omega_{j}\}_{j=1}^{n}$
Algorithm 10 Double Differential Distributional (D3) TD-Learning (Function Approximation)
  Input: a differentiable differential return quantile parameterization: $\hat{\Omega}(s,j,\boldsymbol{w})$, the policy $\pi$ to be used (e.g., $\varepsilon$-greedy), number of per-step reward quantiles, $m$, number of differential return quantiles, $n$, and $\tau$-locations, such that $\tau_{i}=\frac{2i-1}{2m}$ and $\tau_{j}=\frac{2j-1}{2n}$
  Algorithm parameters: step size parameters $\{\alpha$, $\eta_{{}_{\theta}}\}$, quantile Huber loss $\lambda$
  Initialize $\boldsymbol{w}\in\mathbb{R}^{d}$ arbitrarily (e.g. to $\boldsymbol{0}$)
  Initialize per-step reward quantiles $\theta_{1},\theta_{2},\ldots,\theta_{m}$ arbitrarily (e.g. to zero)
  Obtain initial $S$
  while still time to train do
     $A\leftarrow$ action given by $\pi$ for $S$
     Take action $A$, observe $R,S^{\prime}$
     $\bar{R}=\frac{1}{m}\sum_{i=1}^{m}\theta_{i}$
     $\ell=-\sum_{j=1}^{n}\mathbb{E}_{k}\Big{[}h_{\tau_{j}}^{\lambda}\Big{(}R-\bar{R}+\hat{\Omega}(S^{\prime},k,\boldsymbol{w})-\hat{\Omega}(S,j,\boldsymbol{w})\Big{)}\Big{]}$ (See Equation (A.1))
     $\boldsymbol{w}=\boldsymbol{w}+\alpha\frac{\partial\ell}{\partial\boldsymbol{w}}$
     for $i=1,2,\ldots,m$ do
        $\theta_{i}=\theta_{i}+\eta_{{}_{\theta}}\alpha\left(\tau_{i}-\mathds{1}{\{R<\theta_{i}\}}\right)$
     $S=S^{\prime}$
  return $\{\theta_{i}\}_{i=1}^{m}$ and $\boldsymbol{w}$
Algorithm 11 Double Differential Distributional (D3) Q-Learning (Function Approximation)
  Input: a differentiable differential return quantile parameterization: $\hat{\Omega}(s,a,j,\boldsymbol{w})$, the policy $\pi$ to be used (e.g., $\varepsilon$-greedy), number of per-step reward quantiles, $m$, number of differential return quantiles, $n$, and $\tau$-locations, such that $\tau_{i}=\frac{2i-1}{2m}$ and $\tau_{j}=\frac{2j-1}{2n}$
  Algorithm parameters: step size parameters $\{\alpha$, $\eta_{{}_{\theta}}\}$, quantile Huber loss $\lambda$
  Initialize $\boldsymbol{w}\in\mathbb{R}^{d}$ arbitrarily (e.g. to $\boldsymbol{0}$)
  Initialize per-step reward quantiles $\theta_{1},\theta_{2},\ldots,\theta_{m}$ arbitrarily (e.g. to zero)
  Obtain initial $S$
  while still time to train do
     $A\leftarrow$ action given by $\pi$ for $S$
     Take action $A$, observe $R,S^{\prime}$
     $\bar{R}=\frac{1}{m}\sum_{i=1}^{m}\theta_{i}$
     $a^{*}=\text{argmax}_{a^{\prime}}\sum_{j=1}^{n}{\frac{1}{n}\hat{\Omega}(S^{\prime},a^{\prime},j,\boldsymbol{w})}$
     $\ell=-\sum_{j=1}^{n}\mathbb{E}_{k}\Big{[}h_{\tau_{j}}^{\lambda}\Big{(}R-\bar{R}+\hat{\Omega}(S^{\prime},a^{*},k,\boldsymbol{w})-\hat{\Omega}(S,A,j,\boldsymbol{w})\Big{)}\Big{]}$ (See Equation (A.1))
     $\boldsymbol{w}=\boldsymbol{w}+\alpha\frac{\partial\ell}{\partial\boldsymbol{w}}$
     for $i=1,2,\ldots,m$ do
        $\theta_{i}=\theta_{i}+\eta_{{}_{\theta}}\alpha\left(\tau_{i}-\mathds{1}{\{R<\theta_{i}\}}\right)$
     $S=S^{\prime}$
  return $\{\theta_{i}\}_{i=1}^{m}$ and $\boldsymbol{w}$
Algorithm 12 Double Differential Distributional (D3) Actor-Critic
  Input: a differentiable differential return quantile parameterization: $\hat{\Omega}(s,j,\boldsymbol{w})$, a differentiable policy parameterization $\pi(a\mid s,\boldsymbol{u})$, number of per-step reward quantiles, $m$, number of differential return quantiles, $n$, and $\tau$-locations, such that $\tau_{i}=\frac{2i-1}{2m}$ and $\tau_{j}=\frac{2j-1}{2n}$
  Algorithm parameters: step size parameters $\{\alpha$, $\eta_{\pi}$, $\eta_{{}_{\theta}}\}$, quantile Huber loss $\lambda$
  Initialize $\boldsymbol{w}\in\mathbb{R}^{d}$ and $\boldsymbol{u}\in\mathbb{R}^{d^{\prime}}$ (e.g. to $\boldsymbol{0}$)
  Initialize per-step reward quantiles $\theta_{1},\theta_{2},\ldots,\theta_{m}$ arbitrarily (e.g. to zero)
  Obtain initial $S$
  while still time to train do
     $A\sim\pi(\cdot\mid S,\boldsymbol{u})$
     Take action $A$, observe $R,S^{\prime}$
     $\bar{R}=\frac{1}{m}\sum_{i=1}^{m}\theta_{i}$
     $\delta=R-\bar{R}+\mathbb{E}_{j}[\hat{\Omega}(S^{\prime},j,\boldsymbol{w})]-\mathbb{E}_{j}[\hat{\Omega}(S,j,\boldsymbol{w})]$
     $\ell=-\sum_{j=1}^{n}\mathbb{E}_{k}\Big{[}h_{\tau_{j}}^{\lambda}\Big{(}R-\bar{R}+\hat{\Omega}(S^{\prime},k,\boldsymbol{w})-\hat{\Omega}(S,j,\boldsymbol{w})\Big{)}\Big{]}$ (See Equation (A.1))
     $\boldsymbol{w}=\boldsymbol{w}+\alpha\frac{\partial\ell}{\partial\boldsymbol{w}}$
     $\boldsymbol{u}=\boldsymbol{u}+\eta_{\pi}\alpha\delta\nabla\text{ln}\pi(A\mid S,\boldsymbol{u})$
     for $i=1,2,\ldots,m$ do
        $\theta_{i}=\theta_{i}+\eta_{{}_{\theta}}\alpha\left(\tau_{i}-\mathds{1}{\{R<\theta_{i}\}}\right)$
     $S=S^{\prime}$
  return $\{\theta_{i}\}_{i=1}^{m}$ and $\boldsymbol{w}$
Appendix B Convergence Proofs
In this appendix, we present the full convergence proofs for the per-step reward quantile estimates and value function estimates of our tabular Differential Distributional (D2) algorithms (Algorithms 3 and 4). Our general strategy is as follows: we first show that the results from Rowland et al. [2024], which show the almost sure convergence of the discounted return quantile estimates of discounted distributional algorithms, can be used to show the convergence of the per-step reward quantile estimates of our algorithms. We then leverage a two-timescales argument to show the almost sure convergence of the value function estimates. In this regard, Section B.1 contains the proof for the per-step reward quantile estimates, and Section B.2 contains the proof for the value function estimates.
B.1 Convergence of the Per-Step Reward Quantile Estimates
This section contains the proof for the per-step reward quantile estimates. Our general strategy is as follows: we show that the results from Rowland et al. [2024], which show the almost sure convergence of the discounted return quantile estimates of discounted distributional algorithms, can be used to show the almost sure convergence of the per-step reward quantile estimates of our D2 algorithms.
To begin, first consider the discounted return quantile update rule considered in Rowland et al. [2024] (Equation (10) in the main body):
$\Omega^{\gamma}_{i,t+1}=\Omega^{\gamma}_{i,t}+\alpha_{t}\frac{1}{n}\sum_{j=1}^{n}\left[\tau_{i}-\mathds{1}_{\{\psi_{t}<0\}}\right],\;\forall i=1,2,\ldots,n,$ (B.1)
where $\Omega^{\gamma}_{i,t}$ denotes the estimate of the $\tau_{i}$-quantile of the discounted return distribution at time $t$, $\alpha_{t}$ denotes the step size for the update, $\psi_{t}=R_{t}+\gamma\Omega^{\gamma}_{j,t}(S_{t+1})-\Omega^{\gamma}_{i,t}(S_{t})$ (note that Rowland et al. [2024] only considered the prediction case), and $\{\tau_{i}\in(0,1)\}_{i=1}^{n}$, such that $\tau_{i}=\frac{2i-1}{2n},\;i=1,2,\ldots,n$.
To show the convergence of the quantile estimates obtained via the update rule (B.1), Rowland et al. [2024] argued that the update rule (B.1) was a specific instance of the generic update rule:
$\theta_{t+1}=\theta_{t}+\alpha_{t}\left(g(\theta_{t})+M_{t}\right),$ (B.2)
where $M_{t}$ denotes some noise (or error) term, and $g(\theta_{t})\doteq g_{t}\in H(\theta_{t})$, where $H$ denotes a Marchaud map (we will formally define these terms later on). In particular, Rowland et al. [2024] proved the convergence of the generic update rule (B.2), and then showed that the discounted return quantile update rule (B.1) is a specific instance of the generic update rule (B.2), and satisfied the assumptions required for convergence.
Let us now consider the per-step reward quantile update rule from our D2 Algorithms (Equation (13) in the main body):
$\theta_{i,t+1}=\theta_{i,t}+\alpha_{t}\left(\tau_{i}-\mathds{1}_{\{R_{t}<\theta_{i,t}\}}\right),\;\forall i=1,2,\ldots,m,$ (B.3)
where $\theta_{i,t}$ denotes the estimate of the $\tau_{i}$-quantile of the limiting per-step reward distribution, $\alpha_{t}$ denotes the step size for the update, and $R_{t}$ denotes the per-step reward. We claim that the per-step reward quantile update rule (B.3) can also be viewed as an instance of the more generic update rule (B.2). As such, similar to what was done in Rowland et al. [2024] for the discounted return quantile update rule, we will show that the per-step reward quantile update rule (B.3) is indeed an instance of the generic update rule (B.2), and also satisfies the assumptions required for convergence of the more generic update rule (B.2).
We will begin by first summarizing key concepts and relevant results from Rowland et al. [2024]:
Definition B.1.1 (Differential Inclusion).
A differential inclusion is a generalization of an ordinary differential equation (ODE), where the derivative of the unknown function is constrained to lie within a set of possible values, rather than being given by a single-valued function. More formally, given a set-valued map, $H:\mathbb{R}^{m}\rightrightarrows\mathbb{R}^{m}$, the differential inclusion,
$\partial_{t}z_{t}\in H(z_{t}),$
describes a family of trajectories $\{z_{t}\}_{t\geq 0}$ such that the derivative at each time $t$ belongs to the set $H(z_{t})$.
Definition B.1.2 (Definition 11 of Rowland et al. [2024]).
Let $H:\mathbb{R}^{m}\rightrightarrows\mathbb{R}^{m}$ be a set-valued map. The path $\{z_{t}\}_{t\geq 0}$ is a solution to the differential inclusion $\partial_{t}z_{t}\in H(z_{t})$ if there exists an integrable function $g:[0,\infty)\rightarrow\mathbb{R}^{m}$ such that
$z_{t}=\int_{0}^{t}g_{s}\mathrm{d}s$
for all $t\geq 0$, and $g_{t}\in H(z_{t})$ for almost all $t\geq 0$.
Proposition B.1.3 (Proposition 12 of Rowland et al. [2024]).
Consider a set-valued map, $H:\mathbb{R}^{m}\rightrightarrows\mathbb{R}^{m}$, and suppose that $H$ is a Marchaud map; that is,
  • the set $\{(z,h):z\in\mathbb{R}^{m},h\in H(z)\}$ is closed,
  • for all $z\in\mathbb{R}^{m}$, $H(z)$ is non-empty, compact, and convex, and
  • there exists a constant $C>0$ such that for all $z\in\mathbb{R}^{m}$,
    $\max_{h\in H(z)}\|h\|\leq C(1+\|z\|)\,.$
Then, the differential inclusion $\partial_{t}z_{t}\in H(z_{t})$ has a global solution for any initial condition.
Definition B.1.4 (Definition 13 of Rowland et al. [2024]).
Consider a Marchaud map, $H:\mathbb{R}^{m}\rightrightarrows\mathbb{R}^{m}$, and a subset $\Lambda\subseteq\mathbb{R}^{m}$. A continuous function $L:\mathbb{R}^{m}\rightarrow[0,\infty)$ is said to be a Lyapunov function for the differential inclusion $\partial_{t}z_{t}\in H(z_{t})$ and subset $\Lambda$ if for any solution $\{z_{t}\}_{t\geq 0}$ of the differential inclusion and $0\leq s<t$, we have $L(z_{t})<L(z_{s})$ for all $z_{s}\not\in\Lambda$ and $L(z)=0$ for all $z\in\Lambda$.
The above definitions and proposition pertain to the theory of differential inclusions. In essence, the notion of differential inclusions is invoked to account for the fact that the CDF of the limiting per-step reward distribution may not be strictly increasing. That is, there may be discontinuities in the CDF, which means that using a standard ODE framework for the analysis is insufficient as it does not guarantee the existence of a solution. Differential inclusions on the other hand, provide a more general framework in which the dynamics are defined in terms of a set-valued map, $H$, allowing for the solution to take any value within a specified range at any point, including at a given discontinuity. As established in Proposition B.1.3, if the set-valued map, $H$, satisfies the conditions of a Marchaud map, then the existence of global solutions is guaranteed, even if the underlying system exhibits discontinuous behavior.
Having stated the necessary background as it relates to differential inclusions, we are now ready to state the main convergence result from Rowland et al. [2024]:
Theorem B.1.5 (Theorem 14 of Rowland et al. [2024]).
Consider a Marchaud map, $H:\mathbb{R}^{m}\rightrightarrows\mathbb{R}^{m}$, and the corresponding differential inclusion, $\partial_{t}z_{t}\in H(z_{t})$. Suppose there exists a Lyapunov function, $L$, for this differential inclusion and a subset, $\Lambda\subseteq\mathbb{R}^{m}$. Suppose also that we have a sequence $\{\theta_{t}\}_{t\geq 0}$ satisfying
$\theta_{t+1}=\theta_{t}+\alpha_{t}(g(\theta_{t})+M_{t})\,,$
where,
  • $\{\alpha_{t}\}_{t=0}^{\infty}$ satisfy the conditions $\sum_{t=0}^{\infty}\alpha_{t}=\infty$, $\alpha_{t}=o(1/\log(t))$;
  • $g(\theta_{t})\in H(\theta_{t})$ for all $t\geq 0$;
  • $\{M_{t}\}_{t=0}^{\infty}$ is a bounded martingale difference sequence with respect to the natural filtration generated by $\{\theta_{t}\}_{t=0}^{\infty}$. That is, there is an absolute constant $C$ such that $\|M_{t}\|_{\infty}<C$ almost surely, and $\mathbb{E}[M_{t}\mid\theta_{0},\ldots,\theta_{t}]=0$.
If further $\{\theta_{t}\}_{t=0}^{\infty}$ is bounded almost surely (that is, $\sup_{t\geq 0}\|\theta_{t}\|_{\infty}<\infty$ almost surely), then $\theta_{t}\rightarrow\Lambda$ a.s.
Hence, Theorem B.1.5 establishes the convergence of the generic update rule (B.2). As such, as per Theorem B.1.5, in order to establish that the per-step reward quantile estimates generated by the update rule (B.3) converge to the quantiles of the limiting per-step reward distribution, it suffices to verify that the following conditions are satisfied:
  • that the resulting set-valued map, $H$, is a Marchaud map;
  • that the update rule (B.3) can be written in terms of the more generic update rule (B.2), such that $\{M_{t}\}_{t=0}^{\infty}$ is a bounded martingale difference sequence;
  • that $\{\theta_{t}\}_{t=0}^{\infty}$ is bounded almost surely; and
  • that a Lyapunov function, $L$, exists for the associated differential inclusion, $\partial_{t}z_{t}\in H(z_{t})$ and subset $\Lambda$.
We will now show that the above properties hold for the sequence of per-step reward quantile estimates obtained via the update rule (B.3):
Lemma B.1.6 (Marchaud Map for Per-Step Reward Quantile Estimates).
Consider the per-step reward quantile update rule (B.3). Let $F$ denote the CDF of the limiting per-step reward distribution, and let $\tau_{i}\in(0,1)$ denote the $i^{\text{th}}$ target quantile level, with the corresponding quantile value given by $q_{i}=F^{-1}(\tau_{i})$.
Define the set-valued map, $H:\mathbb{R}^{m}\rightrightarrows\mathbb{R}^{m}$, as follows: For any $\theta\in\mathbb{R}^{m}$, let
$H(\theta)_{i}=\tau_{i}-F^{*}(\theta_{i}),$
where,
$F^{*}(\theta_{i})=\begin{cases}\{F(\theta_{i})\},&\text{if }F\text{ is continuous at }\theta_{i}\\ [F(\theta_{i}^{-}),F(\theta_{i})],&\text{if }F\text{ has a jump (or discontinuity) at }\theta_{i}.\end{cases}$
Then $H$ is a Marchaud map, i.e., it satisfies the following properties:
  1. 1.
    The set $\{(\theta,h):\theta\in\mathbb{R}^{m},h\in H(\theta)\}$ is closed;
  2. 2.
    For all $\theta\in\mathbb{R}^{m}$, $H(\theta)$ is non-empty, compact, and convex;
  3. 3.
    There exists a constant $C>0$ such that for all $\theta\in\mathbb{R}^{m}$,
    $\max_{h\in H(\theta)}\|h\|\leq C(1+\|\theta\|).$
Proof.
We verify each condition of a Marchaud map:
Condition 1: To show that the graph of $H$ is closed, consider a convergent sequence $(\theta^{k},h^{k})\rightarrow(\theta,h)$ where $h^{k}\in H(\theta^{k})$ for all $k$. For each component, $i$, we have $h^{k}_{i}=\tau_{i}-p^{k}_{i}$ for some $p^{k}_{i}\in F^{*}(\theta^{k}_{i})$. Since $F^{*}$ takes values in $[0,1]$, the sequence $\{p^{k}_{i}\}$ is bounded and thus has a convergent subsequence with limit $p_{i}$.
If $F$ is continuous at $\theta_{i}$, then by definition of continuity, $F(\theta^{k}_{i})\rightarrow F(\theta_{i})$, and since $p^{k}_{i}=F(\theta^{k}_{i})$ for large enough $k$, it follows that $p_{i}=F(\theta_{i})\in F^{*}(\theta_{i})$.
If $F$ has a jump at $\theta_{i}$, then by the upper hemicontinuity of $F^{*}$, we have $p_{i}\in[F(\theta_{i}^{-}),F(\theta_{i})]=F^{*}(\theta_{i})$.
Therefore, $h_{i}=\tau_{i}-p_{i}$ with $p_{i}\in F^{*}(\theta_{i})$, which implies $h\in H(\theta)$, confirming that the graph of $H$ is closed.
Condition 2: For any $\theta\in\mathbb{R}^{m}$, we have:
  • Non-empty: For each component, $i$, the set $F^{*}(\theta_{i})$ is either the singleton $\{F(\theta_{i})\}$ when $F$ is continuous at $\theta_{i}$, or the closed interval $[F(\theta_{i}^{-}),F(\theta_{i})]$ if $F$ has a jump at $\theta_{i}$. In both cases, $F^{*}(\theta_{i})$ is clearly non-empty, which implies that $H(\theta)_{i}=\tau_{i}-F^{*}(\theta_{i})$ is also non-empty.
  • Compact: Each $F^{*}(\theta_{i})$ is either a singleton $\{F(\theta_{i})\}$ or a closed interval $[F(\theta_{i}^{-}),F(\theta_{i})]$, both of which are compact subsets of $\mathbb{R}$. Since translation (via subtraction by the constant $\tau_{i}$) is a continuous map on $\mathbb{R}$, the image $\tau_{i}-F^{*}(\theta_{i})$ is also compact.
  • Convex: Each $F^{*}(\theta_{i})$ is convex, since both singletons and closed intervals are convex subsets of $\mathbb{R}$. The affine transformation $x\mapsto\tau_{i}-x$ preserves convexity, hence we can conclude that $H(\theta)_{i}=\tau_{i}-F^{*}(\theta_{i})$ is also convex.
Since these properties hold component-wise, $H(\theta)$ as a whole is non-empty, compact, and convex.
Condition 3: For any $\theta\in\mathbb{R}^{m}$ and $h\in H(\theta)$, each component satisfies $h_{i}=\tau_{i}-p_{i}$ for some $p_{i}\in F^{*}(\theta_{i})$. Since $\tau_{i}\in(0,1)$ and $p_{i}\in[0,1]$ (as $F^{*}$ takes values in $[0,1]$), we have $|h_{i}|\leq 1$ for all $i$.
Therefore, when we consider the infinity norm, we have that $\|h\|_{\infty}\leq 1$ for all $h\in H(\theta)$. Hence, we may take $C=1$. This gives:
$\max_{h\in H(\theta)}\|h\|_{\infty}\leq 1\leq C(1+\|\theta\|_{\infty}),$ (B.4)
which satisfies the desired linear growth condition.
As such, given that the three conditions are satisfied, we can conclude that $H$ is a Marchaud map. ∎
Lemma B.1.7 (Bounded Martingale Difference Sequence).
Consider the per-step reward quantile update rule (B.3). This update rule can be re-written in terms of the generic update rule (B.2) as follows:
$\theta_{i,t+1}=\theta_{i,t}+\alpha_{t}(g(\theta_{i,t})+M_{i,t}),\;\forall i=1,2,\ldots,m,$
where,
$g(\theta_{i,t})=\mathbb{E}[\tau_{i}-\mathds{1}_{\{R_{t}<\theta_{i,t}\}}\mid\theta_{i,0},\ldots,\theta_{i,t}],$ (B.5)
and
$M_{i,t}=\tau_{i}-\mathds{1}_{\{R_{t}<\theta_{i,t}\}}-\mathbb{E}[\tau_{i}-\mathds{1}_{\{R_{t}<\theta_{i,t}\}}\mid\theta_{i,0},\ldots,\theta_{i,t}],$ (B.6)
such that the sequence $\{M_{t}\}_{t=0}^{\infty}$ is a bounded martingale difference sequence with respect to the natural filtration generated by $\{\theta_{t}\}_{t=0}^{\infty}$.
Proof.
We need to show that the sequence $\{M_{t}\}_{t\geq 0}$ satisfies two key properties:
1) Martingale Difference: $\mathbb{E}[M_{i,t}\mid\theta_{i,0},\ldots,\theta_{i,t}]=0\;\forall i=1,2,\ldots,m,$ for all $t\geq 0$.
2) Boundedness: There exists a constant $C_{i}>0$, such that $|M_{i,t}|<C_{i}\;\forall i=1,2,\ldots,m,$ a.s. for all $t\geq 0$.
Let us consider an arbitrary $i^{\text{th}}$ component:
Martingale Difference:
Consider Equation (B.6). Taking the conditional expectation with respect to $\mathcal{F}_{i,t}\doteq\theta_{i,0},\ldots,\theta_{i,t}$, we obtain:
$\mathbb{E}[M_{i,t}\mid\mathcal{F}_{i,t}]$ $=\mathbb{E}[\tau_{i}-\mathds{1}_{\{R_{t}<\theta_{i,t}\}}-\mathbb{E}[\tau_{i}-\mathds{1}_{\{R_{t}<\theta_{i,t}\}}\mid\mathcal{F}_{i,t}]\mid\mathcal{F}_{i,t}]$ (B.7)
$=\mathbb{E}[\mathds{1}_{\{R_{t}<\theta_{i,t}\}}\mid\mathcal{F}_{i,t}]-\mathbb{E}[\mathbb{E}[\mathds{1}_{\{R_{t}<\theta_{i,t}\}}\mid\mathcal{F}_{i,t}]\mid\mathcal{F}_{i,t}].$ (B.8)
By the tower property of conditional expectations, we have that:
$\mathbb{E}[\mathbb{E}[\mathds{1}_{\{R_{t}<\theta_{i,t}\}}\mid\mathcal{F}_{i,t}]\mid\mathcal{F}_{i,t}]$ $=\mathbb{E}[\mathds{1}_{\{R_{t}<\theta_{i,t}\}}\mid\mathcal{F}_{i,t}].$ (B.9)
As such,
$\mathbb{E}[M_{i,t}\mid\mathcal{F}_{i,t}]$ $=\mathbb{E}[\mathds{1}_{\{R_{t}<\theta_{i,t}\}}\mid\mathcal{F}_{i,t}]-\mathbb{E}[\mathds{1}_{\{R_{t}<\theta_{i,t}\}}\mid\mathcal{F}_{i,t}]$ (B.10)
$=0,$ (B.11)
which confirms that $\{M_{i,t}\}_{t\geq 0}$ is indeed a martingale difference sequence with respect to the filtration $\{\mathcal{F}_{i,t}\}_{t\geq 0}$.
Boundedness:
Consider Equation (B.6). We can see that the indicator function $\mathds{1}_{\{R_{t}<\theta_{i,t}\}}$ can only take values in $\{0,1\}$. Similarly, its conditional expectation $\mathbb{E}[\mathds{1}_{\{R_{t}<\theta_{i,t}\}}\mid\mathcal{F}_{i,t}]$ is a probability, and hence must lie in the interval $[0,1]$. As such, for all realizations and for all $t\geq 0$:
$|M_{i,t}|$ $=|\tau_{i}-\mathds{1}_{\{R_{t}<\theta_{i,t}\}}-\mathbb{E}[\tau_{i}-\mathds{1}_{\{R_{t}<\theta_{i,t}\}}\mid\mathcal{F}_{i,t}]|$ (B.12)
$\leq\max\{|\mathbb{E}[\mathds{1}_{\{R_{t}<\theta_{i,t}\}}\mid\mathcal{F}_{i,t}]-0|,|\mathbb{E}[\mathds{1}_{\{R_{t}<\theta_{i,t}\}}\mid\mathcal{F}_{i,t}]-1|\}$ (B.13)
$\leq\max\{1,1\}$ (B.14)
$=1$ (B.15)
Taking $C_{i}=1$, we have that $|M_{i,t}|\leq C_{i}$ almost surely for all $t\geq 0$.
Finally, since we chose $i$ arbitrarily, we can extend the above results for all $i=1,2,\ldots,m$. Consequently, we can conclude that the sequence $\{M_{t}\}_{t=0}^{\infty}$ is a bounded martingale difference sequence with respect to the natural filtration generated by $\{\theta_{t}\}_{t=0}^{\infty}$. ∎
Lemma B.1.8 (Boundedness of the Per-Step Reward Quantile Estimates).
Consider the per-step reward quantile update rule (B.3):
$\theta_{i,t+1}=\theta_{i,t}+\alpha_{t}\left(\tau_{i}-\mathds{1}\left\{R_{t}<\theta_{i,t}\right\}\right),\;\forall i=1,2,\ldots,m,$
where $\theta_{i,t}$ denotes the estimate of the $\tau_{i}$-quantile of the limiting per-step reward distribution, $\alpha_{t}$ denotes the step size for the update, and $R_{t}$ denotes the per-step reward.
If:
  1. 1.
    the per-step rewards $\{R_{t}\}_{t\geq 0}$ are bounded (i.e., there exist some finite constants $R_{\text{min}},R_{\text{max}}\in\mathbb{R}$ such that $R_{\text{min}}\leq R_{t}\leq R_{\text{max}}$ for all $t\geq 0$), and
  2. 2.
    the step sizes satisfy $\alpha_{t}>0,\sup_{t\geq 0}\alpha_{t}<\infty$, $\sum_{t=0}^{\infty}\alpha_{t}=\infty$, and $\alpha_{t}=o(1/\log(t))$,
then the sequence $\{\theta_{t}\}_{t=0}^{\infty}$ is bounded almost surely. That is, $\sup_{t\geq 0}|\theta_{i,t}|<\infty$ a.s. for each $i=1,2,\ldots,m$, such that, $\sup_{t\geq 0}\|\theta_{t}\|_{\infty}<\infty$.
Proof.
First, we note that since $\tau_{i}\in(0,1)$, and the indicator function $\mathds{1}\{R_{t}<\theta_{i,t}\}\in\{0,1\}$, the absolute value of the difference $\left|\tau_{i}-\mathds{1}\{R_{t}<\theta_{i,t}\}\right|$ is always at most 1. As such, the magnitude of each per-step update is bounded by the step size. In other words, we have that:
$|\theta_{i,t+1}-\theta_{i,t}|=\alpha_{t}\left|\tau_{i}-\mathds{1}\left\{R_{t}<\theta_{i,t}\right\}\right|\leq\alpha_{t}.$ (B.16)
Moreover, given the step size assumptions, it follows that $\sup_{t\geq 0}\alpha_{t}<\infty$, and hence, that the per-step update magnitudes are uniformly bounded over time.
We will next show that a sequence $\{\theta_{i,t}\}_{t\geq 0}$ for an arbitrary $i^{\text{th}}$ component cannot diverge to either +$\infty$+ ∞ or $\infty$- ∞:
Upper Bound:
We proceed by contradiction. Suppose there exists a sample path where $\theta_{i,t}\rightarrow+\infty$ as $t\rightarrow\infty$. Then there must exist some time $T$ such that $\theta_{i,t}>R_{\text{max}}$ for all $t\geq T$.
Consider the update rule for this sample path. For all $t\geq T$, we have that:
$\theta_{i,t+1}-\theta_{i,t}$ $=\alpha_{t}(\tau_{i}-\mathds{1}\left\{R_{t}<\theta_{i,t}\right\})$ (B.17a)
$=\alpha_{t}(\tau_{i}-1)$ (B.17b)
$=-\alpha_{t}(1-\tau_{i}).$ (B.17c)
Since $\tau_{i}\in(0,1)$, we have that $-\alpha_{t}(1-\tau_{i})<0$, which means that the change in $\theta_{i,t}$ is negative.
Now consider the cumulative effect over time $T$ through some time $T+k-1$:
$\theta_{i,T+k}-\theta_{i,T}$ $=\sum_{t=T}^{T+k-1}(\theta_{i,t+1}-\theta_{i,t})$ (B.18a)
$=\sum_{t=T}^{T+k-1}-\alpha_{t}(1-\tau_{i})$ (B.18b)
$=-(1-\tau_{i})\sum_{t=T}^{T+k-1}\alpha_{t}.$ (B.18c)
Looking at Equation (B.18c), we know by assumption that $\sum_{t=0}^{\infty}\alpha_{t}=\infty$. As such, there must exist some timestep, $k^{*}$, such that eventually, the sample path will eventually return to and cross $R_{\text{max}}$. In other words, we will eventually have:
$\theta_{i,T}-(1-\tau_{i})\sum_{t=T}^{T+k^{*}-1}\alpha_{t}\leq R_{\text{max}},$ (B.19)
and thus, that
$\theta_{i,T+k^{*}}=\theta_{i,T}-(1-\tau_{i})\sum_{t=T}^{T+k^{*}-1}\alpha_{t}\leq R_{\text{max}},$ (B.20)
which contradicts the assumption that $\theta_{i,t}>R_{\text{max}}$ for all $t\geq T$.
Lower Bound:
We proceed by contradiction. Suppose there exists a sample path where $\theta_{i,t}\rightarrow-\infty$ as $t\rightarrow\infty$. Then there must exist some time $T$ such that $\theta_{i,t}<R_{\text{min}}$ for all $t\geq T$.
Consider the update rule for this sample path. For all $t\geq T$, we have that:
$\theta_{i,t+1}-\theta_{i,t}$ $=\alpha_{t}(\tau_{i}-\mathds{1}\left\{R_{t}<\theta_{i,t}\right\})$ (B.21a)
$=\alpha_{t}\tau_{i}.$ (B.21b)
Since $\tau_{i}>0$, the change in $\theta_{i,t}$ is positive.
Now consider the cumulative effect over time $T$ through some time $T+k-1$:
$\theta_{i,T+k}-\theta_{i,T}=\tau_{i}\sum_{t=T}^{T+k-1}\alpha_{t}.$ (B.22)
Looking at Equation (B.22), we know by assumption that $\sum_{t=0}^{\infty}\alpha_{t}=\infty$. As such, there must exist some timestep, $k^{*}$, such that eventually, the sample path will eventually return to and cross $R_{\text{min}}$. In other words, we will eventually have:
$\theta_{i,T}+\tau_{i}\sum_{t=T}^{T+k^{*}-1}\alpha_{t}\geq R_{\text{min}},$ (B.23)
and thus, that
$\theta_{i,T+k^{*}}=\theta_{i,T}+\tau_{i}\sum_{t=T}^{T+k^{*}-1}\alpha_{t}\geq R_{\text{min}},$ (B.24)
which contradicts the assumption $\theta_{i,t}<R_{\text{min}}$ for all $t\geq T$.
As such, we have shown that $\theta_{i,t}$ cannot diverge to +$\infty$+ ∞ or $\infty$- ∞ along any sample path. Hence, the sequence $\{\theta_{i,t}\}$ must remain within a bounded interval almost surely. Moreover, because we chose $i$ arbitrarily, we can extend the above results for all $i=1,2,\ldots,m$. Consequently, we can conclude that $\sup_{t\geq 0}\|\theta_{t}\|_{\infty}<\infty$. ∎
Lemma B.1.9 (Existence of a Lyapunov Function for the Per-Step Reward Quantile Estimates).
Consider the differential inclusion $\partial_{t}z_{t}\in H(z_{t})$ associated with the per-step reward quantile update rule (B.3), where $H$ is the Marchaud map defined in Lemma B.1.6. Let $F$ be the CDF of the limiting per-step reward distribution, and let $q_{i}=F^{-1}(\tau_{i})$ denote the true $\tau_{i}$-quantile of this distribution.
Define the subset $\Lambda=\{\theta\in\mathbb{R}^{m}:\theta_{i}\in q_{i}\;\;\text{for all }i=1,2,\dots,m\}$, where:
$q_{i}=\begin{cases}F^{-1}(\tau_{i}),&\text{if }F^{-1}(\tau_{i})\text{ is unique},\\ [a_{i},b_{i}],&\text{if }F\text{ has a flat region at level }\tau_{i},\end{cases}$
with $a_{i}=\inf\{x:F(x)\geq\tau_{i}\}$ and $b_{i}=\sup\{x:F(x)\leq\tau_{i}\}$.
The function $L:\mathbb{R}^{m}\rightarrow[0,\infty)$ defined by:
[…]
where the integral is defined as:
[…]
is a Lyapunov function for the differential inclusion $\partial_{t}z_{t}\in H(z_{t})$ and the subset $\Lambda$.
Proof.
To establish that $L$ is a Lyapunov function, we must show that:
  1. 1.
    $L$ is continuous;
  2. 2.
    $L(\theta)=0$ for all $\theta\in\Lambda$; and
  3. 3.
    For any solution $\{z_{t}\}_{t\geq 0}$ of the differential inclusion and $0\leq s<t$, we have $L(z_{t})<L(z_{s})$ for all $z_{s}\not\in\Lambda$.
Continuity of $L$:
First, let us properly define the integral in the definition of $L$. For each component, $i$, we define:
[…]
Since $F$ is a CDF, it is right-continuous by definition. Thus, for any discontinuity in $F$, we interpret $F(s)$ as its right-hand limit, thereby ensuring that the Riemann-Stieltjes integral remains well-defined. Moreover, the function $F(s)-\tau_{i}$ is bounded (since $\tau_{i}\in(0,1)$ and $F(s)\in[0,1]$) and only has jump discontinuities (due to the properties of $F$), making the integral well-defined for any choice of $\theta_{i}$ and $q_{i}$.
Now, for the continuity of $L$, consider a sequence $\{\theta^{k}\}$ converging to $\theta$. For each component, $i$, we claim that:
[…]
This convergence holds because, as discussed in the above paragraph, the integrand $s\mapsto F(s)-\tau_{i}$ is bounded and has at most countably many discontinuities. More specifically, it is well-known that for such functions, the integral is continuous with respect to its upper limit of integration. That is, if $\theta_{i}^{k}\to\theta_{i}$, then the corresponding integrals converge. Since this holds for each $i$-component, and $L(\theta)$ is a finite sum over these integrals, it follows that $L(\theta^{k})\to L(\theta)$. Therefore, $L$ is continuous on $\mathbb{R}^{m}$.
$L(\theta)=0$ for all $\theta\in\Lambda$:
Let $\theta\in\Lambda$. By definition, this means that $\theta_{i}\in q_{i}$ for all $i$. There are two cases that we need to consider:
Case 1: If $F^{-1}(\tau_{i})$ is unique (i.e., $q_{i}$ is a singleton), then $\theta_{i}=q_{i}$, which implies:
[…]
Case 2: If $F$ has a flat region at level $\tau_{i}$ (i.e., $q_{i}=[a_{i},b_{i}]$), then $\theta_{i}\in[a_{i},b_{i}]$, and for all $s\in[a_{i},b_{i}]$, we have $F(s)=\tau_{i}$, so:
[…]
Thus, each component contributes zero to the sum, and $L(\theta)=0$ for all $\theta\in\Lambda$.
$L(z_{t})<L(z_{s})$ for all $z_{s}\not\in\Lambda$ and $0\leq s<t$:
Let $\{z_{t}\}_{t\geq 0}$ be a solution to $\partial_{t}z_{t}\in H(z_{t})$, and let $0\leq s<t$. From Definition B.1.2, there exists an integrable function $g:[0,\infty)\rightarrow\mathbb{R}^{m}$, such that:
[…]
where $g_{u}\in H(z_{u})$ for almost all $u\geq 0$.
By the chain rule and the fundamental theorem of calculus, for almost every $t$, we have that:
$\frac{d}{dt}L(z_{t})=\sum_{i=1}^{m}\frac{\partial L}{\partial z_{i,t}}(z_{t})\cdot\frac{dz_{i,t}}{dt}=\sum_{i=1}^{m}(F(z_{i,t})-\tau_{i})\cdot g_{i,t},$ (B.30)
where $g_{i,t}$ is the $i$-th component of $g_{t}$.
Now, by definition of the Marchaud map, $H$, we have that $g_{i,t}=\tau_{i}-p_{i,t}$, where $p_{i,t}\in F^{*}(z_{i,t})$ (see Lemma B.1.6). Hence:
$\frac{d}{dt}L(z_{t})=\sum_{i=1}^{m}(F(z_{i,t})-\tau_{i})\cdot(\tau_{i}-p_{i,t}).$ (B.31)
We will now analyze the sign of the term $(F(z_{i,t})-\tau_{i})\cdot(\tau_{i}-p_{i,t})$ for each component, $i$, when $z_{i,t}\notin q_{i}$. There are two cases to consider:
Case 1: $z_{i,t}>\sup q_{i}$: Since $q_{i}=F^{-1}(\tau_{i})$ represents the set of all values where $F$ takes the value $\tau_{i}$, and $F$ is non-decreasing, we have that $F(z_{i,t})>\tau_{i}$ for any $z_{i,t}>\sup q_{i}$, such that $F(z_{i,t})-\tau_{i}>0$. Now consider the $\tau_{i}-p_{i,t}$ term. By definition of the Marchaud map, $H$, when $F$ is continuous at $z_{i,t}$, we have that $p_{i,t}=F(z_{i,t})$. Now, since $z_{i,t}>\sup q_{i}$ and $F$ is non-decreasing, we know that $F(z_{i,t}^{-})>F(\sup q_{i})=\tau_{i}$. Similarly, when $F$ has a jump discontinuity at $z_{i,t}$, the value $p_{i,t}$ may lie anywhere in the interval $[F(z_{i,t}^{-}),F(z_{i,t})]$, where $F(z_{i,t}^{-})$ denotes the left-hand limit of $F$ at $z_{i,t}$. In this case, since $F(z_{i,t}^{-})>\tau_{i}$, we know that $p_{i,t}\geq F(z_{i,t}^{-})>\tau_{i}$. As such, in either case we have that $\tau_{i}-p_{i,t}<0$. Hence, we can conclude that $(F(z_{i,t})-\tau_{i})\cdot(\tau_{i}-p_{i,t})<0$.
Case 2: $z_{i,t}<\inf q_{i}$: Since $q_{i}=F^{-1}(\tau_{i})$ represents the set of all values where $F$ takes the value $\tau_{i}$, and $F$ is non-decreasing, we have that $F(z_{i,t})<\tau_{i}$ for any $z_{i,t}<\inf q_{i}$, such that $F(z_{i,t})-\tau_{i}<0$. Now consider the $\tau_{i}-p_{i,t}$ term. By definition of the Marchaud map, $H$, when $F$ is continuous at $z_{i,t}$, we have $p_{i,t}=F(z_{i,t})$. Now, since $z_{i,t}<\inf q_{i}$ and $F$ is non-decreasing, we know that $F(z_{i,t})<F(\inf q_{i}^{-})\leq\tau_{i}$, where $F(\inf q_{i}^{-})$ is the left-hand limit of $F$ at $\inf q_{i}$. Similarly, when $F$ has a jump discontinuity at $z_{i,t}$, the value $p_{i,t}$ may take any value in the interval $[F(z_{i,t}^{-}),F(z_{i,t})]$, where $F(z_{i,t}^{-})$ denotes the left-hand limit of $F$ at $z_{i,t}$. In this case, since $F(z_{i,t})<\tau_{i}$, we know that $p_{i,t}\leq F(z_{i,t})<\tau_{i}$. As such, in either case we have that $\tau_{i}-p_{i,t}>0$. Hence, we can conclude that $(F(z_{i,t})-\tau_{i})\cdot(\tau_{i}-p_{i,t})<0$.
As such, in both cases we can conclude that $(F(z_{i,t})-\tau_{i})\cdot(\tau_{i}-p_{i,t})<0$.
Now, let us define $D(z_{t})=\{i\in\{1,\ldots,m\}:z_{i,t}\not\in q_{i}\}$. By the above analysis, for each $i\in D(z_{t})$, the corresponding term in the sum is strictly negative. Hence, if $z_{t}\not\in\Lambda$, then $D(z_{t})$ is non-empty, and we have:
$\frac{d}{dt}L(z_{t})=\sum_{i=1}^{m}(F(z_{i,t})-\tau_{i})\cdot(\tau_{i}-p_{i,t})=\sum_{i\in D(z_{t})}(F(z_{i,t})-\tau_{i})\cdot(\tau_{i}-p_{i,t})<0.$
Therefore, if $z_{t}\not\in\Lambda$, we can conclude that $\frac{d}{dt}L(z_{t})<0$.
As such, integrating over $[s,t]$ gives:
[…]
which implies that $L(z_{t})<L(z_{s})$ for all $z_{s}\not\in\Lambda$ and $0\leq s<t$.
As such, we have now verified the three conditions, and can therefore conclude that $L$ is a Lyapunov function for $\partial_{t}z_{t}\in H(z_{t})$ and $\Lambda$. ∎
With the above Lemmas, we have now established that the necessary conditions for convergence hold for the per-step reward quantile update rule (B.3). In particular, we are now in a position to apply Theorem B.1.5, which allows us to conclude that the quantile estimates generated by the per-step reward update rule (B.3) converge, almost surely, to the quantiles of the limiting per-step reward distribution. We formalize this argument as Theorem B.1.10 below:
Theorem B.1.10 (Convergence of Per-Step Reward Quantile Estimates).
Consider the per-step reward quantile update rule (B.3):
$\theta_{i,t+1}=\theta_{i,t}+\alpha_{t}\left(\tau_{i}-\mathds{1}\{R_{t}<\theta_{i,t}\}\right),\quad\forall i=1,2,\ldots,m,$
where $\theta_{i,t}$ is the estimate of the $\tau_{i}$-quantile of the limiting per-step reward distribution, $\alpha_{t}$ is the step size, and $R_{t}\in\mathcal{R}$ is the per-step reward.
If:
  1. 1.
    the per-step rewards $\{R_{t}\}_{t\geq 0}$ are bounded,
  2. 2.
    the step sizes $\{\alpha_{t}\}_{t\geq 0}$ satisfy $\alpha_{t}>0,$, $\sup_{t\geq 0}\alpha_{t}<\infty$, $\sum_{t=0}^{\infty}\alpha_{t}=\infty$, and $\alpha_{t}=o(1/\log(t))$, and
  3. 3.
    the distribution of per-step rewards converges to a unique limiting distribution with CDF, $F$ (e.g. via Assumptions 4.2 and 4.3),
then, the sequence $\{\theta_{i,t}\}_{t\geq 0}$ converges, almost surely, to the set $q_{i},\;\forall i=1,2,\ldots,m$, where:
$q_{i}=\begin{cases}F^{-1}(\tau_{i}),&\text{if }F^{-1}(\tau_{i})\text{ is unique},\\ [a_{i},b_{i}],&\text{if }F\text{ has a flat region at level }\tau_{i},\end{cases}$
with $a_{i}=\inf\{x:F(x)\geq\tau_{i}\}$ and $b_{i}=\sup\{x:F(x)\leq\tau_{i}\}$, as $t\rightarrow\infty$.
Proof.
In Lemma B.1.7, we showed that the per-step reward quantile update rule (B.3) can be written in terms of the more generic stochastic approximation update rule (B.2). Moreover, in Lemmas B.1.6, B.1.7, B.1.8, and B.1.9, we showed that:
  • The resulting set-valued map, $H$, is a Marchaud map;
  • The resulting noise sequence, $\{M_{t}\}_{t\geq 0}$, is a bounded martingale difference sequence;
  • The sequence $\{\theta_{i,t}\}_{t\geq 0}$ is bounded almost surely $\forall i=1,2,\ldots,m$; and
  • There exists an appropriate Lyapunov function $L$ for the resulting differential inclusion, $\partial_{t}z_{t}\in H(z_{t})$, and the subset, $\Lambda\doteq\{\theta\in\mathbb{R}^{m}:\theta_{i}\in q_{i}\;\forall i=1,2,\dots,m\}$.
As such, these properties collectively satisfy all of the conditions of Theorem B.1.5 (i.e., Theorem 14 of Rowland et al. [2024]), which guarantees the almost sure convergence of updates rules that are of the form of the generic update rule (B.2). Therefore, in the case of the per-step reward quantile update rule (B.3), we can conclude that $\theta_{i,t}\rightarrow q_{i}$, almost surely, for each $i=1,2,\dots,m$.
This completes the proof. ∎
B.2 Convergence of the Value Function Estimates
This section contains the convergence proofs for the value function estimates of the tabular D2 TD-learning and tabular D2 Q-learning algorithms. Our general strategy is as follows: we build upon existing results from Section B.1, in combination with a two-timescales argument, to show the almost sure convergence of the value function estimates of our algorithms.
In this section, we adopt the following notation:
  • For a given vector $x$, let $\sum x$ denote the sum of all elements in $x$, such that $\sum x\doteq\sum_{i}x(i)$.
  • Let $e$ denote an all-ones vector.
B.2.1 Convergence Proof for the Tabular D2 TD-learning Value Function Estimates
In this section, we present the proof for the convergence of the value function estimates of the D2 TD-learning algorithm (Algorithm 3).
To begin, consider an MDP, $\mathcal{M}\doteq\langle\mathcal{S},\mathcal{A},\mathcal{R},p\rangle$, with policy, $\pi$. Given a state $s\in\mathcal{S}$ and discrete step $n\geq 0$, let $A_{n}(s)\sim\pi(\cdot\mid s)$ denote the action selected, let $R_{n}(s,A_{n}(s))\in\mathcal{R}$ denote a sample of the resulting reward, and let $S^{\prime}_{n}(s,A_{n}(s))\sim p(\cdot,\cdot\mid s,a)$ denote a sample of the resulting state. Let $\{Y_{n}\}$ be a set-valued process taking values in the set of nonempty subsets of $\mathcal{S}$, such that: $Y_{n}=\{s:s$ component of the $|\mathcal{S}|$-sized table of state-value estimates, $V$, that was updated at step $n\}$. Let $\nu(n,s)\doteq\sum_{j=0}^{n}I\{s\in Y_{j}\}$, where $I$ is the indicator function, such that $\nu(n,s)$ represents the number of times that $V(s)$ was updated up until step $n$.
Now consider the following update rule for $n\geq 0$:
$V_{n+1}(s)$ $\doteq V_{n}(s)+\alpha_{\nu(n,s)}\delta_{n}(s)I\{s\in Y_{n}\},\quad\forall s\in\mathcal{S},$ (B.32)
where,
$\begin{split}\delta_{n}(s)&\doteq R_{n}(s,A_{n}(s))-\bar{R}_{n}+V_{n}(S_{n}^{\prime}(s,A_{n}(s)))-V_{n}(s).\end{split}$ (B.33)
Here, $\bar{R}_{n}$ denotes the estimate of the average-reward induced by following policy $\pi$, $\bar{r}_{\pi}$ (see Equation (2)), $\delta_{n}(s)$ denotes the TD error, and $\alpha_{\nu(n,s)}$ denotes the step size at time step $n$ for state $s$.
We will now show that the value function update rule for the D2 TD-learning algorithm is a special case of the update rule (B.32). To this end, consider a sequence of experience from our MDP $\mathcal{M}$: $S_{t},A_{t}(S_{t}),R_{t+1},S_{t+1},\ldots$. Now recall the set-valued process $\{Y_{n}\}$. If we let $n$ = time step $t$, we have:
$Y_{t}(s)=\begin{cases}1,s=S_{t},\\ 0,\text{ otherwise,}\end{cases}$
as well as $S^{\prime}_{n}(S_{t},A_{t}(S_{t}))=S_{t+1}$, and $R_{n}(S_{t},A_{t})=R_{t+1}$.
Hence, the update rule (B.32) becomes:
$V_{t+1}(S_{t})$ $\doteq V_{t}(S_{t})+\alpha_{\nu(t,S_{t})}\delta_{t}\text{\; and \; }V_{t+1}(s)\doteq V_{t}(s),\forall s\neq S_{t},$ (B.34)
which is D2 TD-learning’s value function update rule with $\alpha_{\nu(t,S_{t})}$ denoting the step size at time $t$.
As such, if we can show that the update rule (B.32) converges, then it would imply that the value function update rule for the D2 TD-learning algorithm also converges. To this end, we will now specify the assumptions that are needed to ensure convergence:
Assumption B.2.1 (Unichain Assumption).
The Markov chain induced by the policy is unichain. That is, the induced Markov chain consists of a single recurrent class and a potentially-empty set of transient states.
Assumption B.2.2 (Step Size Assumption).
$\alpha_{n}>0$, $\sup_{n\geq 0}\alpha_{n}<\infty$, $\sum_{n=0}^{\infty}\alpha_{n}=\infty$, $\sum_{n=0}^{\infty}\alpha_{n}^{2}<\infty$, and $\alpha_{n}=o(1/\log(n))$.
Assumption B.2.3 (Asynchronous Step Size Assumption 1).
Let []delimited-[][$\cdot$][ ⋅ ] denote the integer part of ()($\cdot$)( ⋅ ). For $x\in(0,1)$,
$\sup_{i}\frac{\alpha_{[xi]}}{\alpha_{i}}<\infty$
and
$\frac{\sum_{j=0}^{[yi]}\alpha_{j}}{\sum_{j=0}^{i}\alpha_{j}}\to 1$
uniformly in $y\in[x,1]$.
Assumption B.2.4 (Asynchronous Step Size Assumption 2).
There exists $\Delta>0$ such that
$\liminf_{n\to\infty}\frac{\nu(n,s)}{n+1}\geq\Delta,$
a.s., for all $s\in\mathcal{S}$.
Furthermore, for all $x>0$, and
$N(n,x)=\min\Bigg{\{}m\geq n:\sum_{i=n+1}^{m}\alpha_{i}\geq x\Bigg{\}},$
the limit
$\lim_{n\to\infty}\frac{\sum_{i=\nu(n,s)}^{\nu(N(n,x),s)}\alpha_{i}}{\sum_{i=\nu(n,s^{\prime})}^{\nu(N(n,x),s^{\prime})}\alpha_{i}}$
exists a.s. for all $s,s^{\prime}$.
Assumption B.2.5 (Quantile Step Size Assumption).
The quantile step sizes, $\{\eta_{{}_{\theta}}\alpha_{n}\}_{n\geq 0}$, satisfy the following properties: $\eta_{{}_{\theta}}\alpha_{n}>0$, $\sup_{n\geq 0}\eta_{{}_{\theta}}\alpha_{n}<\infty$, $\sum_{n=0}^{\infty}\eta_{{}_{\theta}}\alpha_{n}=\infty$, $\sum_{n=0}^{\infty}(\alpha_{n}^{2}+\eta_{{}_{\theta}}^{2}\alpha_{n}^{2})<\infty$, and $\alpha_{n}/(\eta_{{}_{\theta}}\alpha_{n})\to 0$.
Assumptions B.2.2, B.2.3, and B.2.4, most of which originate from Borkar [1998], outline the step size requirements needed to show the convergence of stochastic approximation algorithms. Assumptions B.2.2 and B.2.3 can be satisfied with step size sequences that decrease to 0 appropriately, such as $1/n^{p}$ with $0.5<p\leq 1$. Assumption B.2.4 first requires that the limiting ratio of visits to any given state, compared to the total number of visits to all states, is greater than or equal to some fixed positive value. The assumption then requires that the relative update frequency between any two states is finite. Assumption B.2.5 ensures that the quantile step size decreases to 0 appropriately.
We are now ready to state the convergence theorem:
Theorem B.2.6 (Convergence of D2 TD-learning Value Function Estimates).
If Assumptions B.2.1B.2.5 hold, then the update rule (B.32) converges, almost surely, $V_{n}\to v_{\infty}\text{ as }n\to\infty$, thereby implying that the D2 TD-learning value function estimates (B.34) converge almost surely to $v_{\infty}\text{ as }n\to\infty\textbf{}$, where $v_{\infty}$ denotes the solution to the Bellman equation (4).
We prove this theorem in the following section. To do so, use a two-timescales argument, such that we leverage Theorem 2 in Section 6 of Borkar [2009], along with the results from Theorems 4.5 and B.1.10.
B.2.2 Proof of Theorem B.2.6
To begin, let us consider Assumption B.2.5. In particular, $\alpha_{n}/(\eta_{{}_{\theta}}\alpha_{n})\to 0$ implies that the quantile step size, $\eta_{{}_{\theta}}\alpha_{n}$, decreases to 0 at a slower rate than the value function step size, $\alpha_{n}$. This implies that the quantile updates move on a faster timescale compared to the value function update. Hence, as argued in Section 6 of Borkar [2009], the (faster) quantile updates (B.3) view the (slower) value function update (B.32) as quasi-static, while the (slower) value function update views the (faster) quantile updates as nearly equilibrated (as we will discuss below, the existence of such an equilibrium point was shown in Theorem B.1.10).
Importantly, given the results of Lemma 4.4, Theorem B.1.10, and Theorem 4.5, the two-timescales argument allows us to view the average-reward estimate, $\bar{R}_{n}$, as nearly equilibrated at $\bar{r}_{\pi}$. In particular, given Lemma 4.4, we have that $\bar{R}_{n}\doteq\frac{1}{m}\sum_{i=1}^{m}\theta_{i,n}$, where $m$ denotes the number of quantiles, and $\theta_{i,n}$ denotes the $i^{\text{th}}$ quantile estimate at step $n$. Now, given Theorem B.1.10, we now that the quantile estimates converge to the desired quantiles of the limiting per-step reward distribution. As such, by Theorem 4.5, we have that $\bar{R}_{n}\to\bar{r}_{\pi}$. Consequently, by the two-timescales argument, the value function update views the average-reward estimate, $\bar{R}_{n}$, as nearly equilibrated at $\bar{r}_{\pi}$ (that is, we can essentially treat $\bar{r}_{\pi}$ as a known constant).
We now proceed to show the convergence of the value function estimates. To this end, consider the asynchronous value function update (B.32). This update is of the same form as Equation 7.1.2 of Borkar [2009]. As such, to show the convergence of the value function estimates, we can apply the result in Section 7.4 of Borkar [2009], which shows the convergence of asynchronous updates of the same form as Equation 7.1.2. To apply this result, given Assumptions B.2.3 and B.2.4, we only need to show the convergence of the synchronous version of the value function update:
$V_{n+1}(s)=V_{n}(s)+\alpha_{n}\left(h(V_{n})+M_{n+1}\right),\quad\forall s\in\mathcal{S},$ (B.35)
where,
$h(V_{n})(s)$ $\doteq\sum_{s^{\prime},r}p(s^{\prime},r\mid s,a)(r+V_{n}(s^{\prime}))-V_{n}(s)-\bar{r}_{\pi},$
$=T(V_{n})(s)-V_{n}(s)-\bar{r}_{\pi},\text{ and}$
$M_{n+1}(s)$ $\doteq R_{n}(s,A_{n}(s))+V_{n}(S_{n}^{\prime}(s,A_{n}(s)))-V_{n}(s)-\bar{r}_{\pi}-h(V_{n})(s).$
To show the convergence of the synchronous update (B.35) under the two-timescales argument, we can apply the result of Theorem 2 in Section 6 of Borkar [2009] to show that $V_{n}\to v_{\infty}$ a.s. as $n\to\infty$. This theorem requires that 3 assumptions be satisfied. As such, we will now show, via Lemmas B.2.7 - B.2.9, that these 3 assumptions are indeed satisfied.
Lemma B.2.7.
The quantile update rules, which can be written in the following form: $\theta_{i,n+1}=\theta_{i,n}+\eta_{{}_{\theta}}\alpha_{n}(g(\theta_{i,n})+M_{i,n})$ (see Lemma B.1.7), each have a globally asymptotically stable equilibrium, $\theta_{i},\;\forall i=1,2,\ldots m$, where $m$ denotes the number of quantiles.
Proof.
This was shown in Theorem B.1.10. ∎
Lemma B.2.8.
The value function update rule (B.35) has a globally asymptotically stable equilibrium, $v_{\infty}$.
Proof.
The value function update rule (B.35) is of the same form as the update rule considered in Lemmas B.6 and B.7 of Wan et al. [2021] (with the importance sampling ratio set to $1.0$). Given that we adhere to all the assumptions required for the aforementioned Lemmas from Wan et al. [2021], we can leverage their results, which show that the value function update rule (B.35) has a globally asymptotically stable equilibrium, $v_{\infty}$. ∎
Lemma B.2.9.
$\sup_{n}(||V_{n}||+\sum_{i=1}^{m}{||\theta_{i,n}||)}<\infty$ a.s.
Proof.
It was shown in Lemma B.1.8 that $\sup_{n}(||\theta_{i,n}||)<\infty\;\forall i=1,2,\ldots,m$ a.s. Hence, we only need to show that $\sup_{n}(||V_{n}||)<\infty$ a.s. To this end, we can apply Theorem 7 in Section 3 of Borkar [2009]. This theorem requires 4 assumptions:
  • (A1) The function $h$ is Lipschitz. That is, $||h(x)-h(y)||\leq L||x-y||$ for some $0<L<\infty$.
  • (A2) The sequence $\{\alpha_{n}\}$ satisfies $\alpha_{n}>0$, $\sum\alpha_{n}=\infty$, and $\sum\alpha_{n}^{2}<\infty$.
  • (A3) $\{M_{n}\}$ is a martingale difference sequence that is square-integrable.
  • (A4) The functions $h_{d}(x)\doteq h(dx)/d$, $d\geq 1,x\in\mathbb{R}^{k}$, satisfy $h_{d}(x)\to h_{\infty}(x)$ as $d\to\infty$, uniformly on compacts for some $h_{\infty}\in C(\mathbb{R}^{k})$. Furthermore, the ODE $\dot{x}_{t}=h_{\infty}(x_{t})$ has the origin as its unique globally asymptotically stable equilibrium.
We first note that Assumption (A1) is satisfied given that both the $\sum$ and $T$ operators are Lipschitz. Moreover, we note that Assumption B.2.2 satisfies Assumption (A2).
We now show that Assumption (A3) is satisfied:
Let $\mathcal{F}_{n}\doteq\sigma(V_{i},M_{i},i\leq n),n\geq 0$ denote an increasing family of $\sigma$-fields. We have that, for any $s\in\mathcal{S}$:
$\mathbb{E}[M_{n+1}(s)\mid\mathcal{F}_{n}]$ $=\mathbb{E}\left[R_{n}(s,A_{n}(s))+V_{n}(S_{n}^{\prime}(s,A_{n}(s)))-V_{n}(s)-\bar{r}_{\pi}-h(V_{n})(s)\mid\mathcal{F}_{n}\right]$ (B.36)
$=\mathbb{E}\left[R_{n}(s,A_{n}(s))+V_{n}(S_{n}^{\prime}(s,A_{n}(s)))-V_{n}(s)-\bar{r}_{\pi}\mid\mathcal{F}_{n}\right]-h(V_{n})(s)$ (B.37)
$=\mathbb{E}\left[R_{n}(s,A_{n}(s))+V_{n}(S_{n}^{\prime}(s,A_{n}(s)))\mid\mathcal{F}_{n}\right]-V_{n}(s)-\bar{r}_{\pi}-h(V_{n})(s)$ (B.38)
$=\mathbb{E}[R_{n}(s,A_{n}(s))+V_{n}(S_{n}^{\prime}(s,A_{n}(s)))\mid\mathcal{F}_{n}]-T(V_{n})(s)$ (B.39)
$=0.$ (B.40)
Moreover, it can be easily shown via the triangle inequality that $\mathbb{E}[||M_{n+1}||^{2}\mid\mathcal{F}_{n}]\leq K(1+||V_{n}||^{2})$ for some constant $K>0$. As such, Assumption (A3) is satisfied.
Finally, we show that Assumption (A4) is also satisfied:
We first note that:
$h_{\infty}(x)=\lim_{d\to\infty}h_{d}(x)=\lim_{d\to\infty}\frac{T(dx)-dx-\bar{r}_{\pi}e}{d}=T_{0}(x)-x,$ (B.41)
where
$T_{0}(x)\doteq\sum_{a}\pi(a\mid s)\sum_{s^{\prime},r}p(s^{\prime},r\mid s,a)x(s^{\prime}).$ (B.42)
Clearly, the function $h_{\infty}$ is continuous in every $x\in\mathbb{R}^{k}$. As such, we have that $h_{\infty}\in C(\mathbb{R}^{k})$.
Next, consider the ODE $\dot{x}_{t}=h_{\infty}(x_{t})=T_{0}(x_{t})-x_{t}$. This ODE has the origin as an equilibrium since $T_{0}(0)-0=0$. Moreover, we can again leverage the results of Lemmas B.6 and B.7 from Wan et al. [2021] to conclude that the origin is the unique globally asymptotically stable equilibrium for this ODE. Hence, Assumption A4 is satisfied.
As such, Assumptions (A1) - (A4) are all verified, meaning that we can apply the results of Theorem 7 in Section 3 of Borkar [2009] to conclude that $\sup_{n}(||V_{n}||)<\infty$ a.s., and hence, that $\sup_{n}(||V_{n}||+\sum_{i=1}^{m}{||\theta_{i,n}||)}<\infty$ a.s. ∎
As such, we have now verified the 3 assumptions required by Theorem 2 in Section 6 of Borkar [2009], which means that we can apply the result of the theorem to conclude that $V_{n}\to v_{\infty}$ a.s. as $n\to\infty$.
This completes the proof of Theorem B.2.6.
B.2.3 Convergence Proof for the Tabular D2 Q-learning Value Function Estimates
In this section, we present the proof for the convergence of the value function estimates of the D2 Q-learning algorithm (Algorithm 4).
To begin, consider an MDP $\mathcal{M}\doteq\langle\mathcal{S},\mathcal{A},\mathcal{R},p\rangle$. Given a state $s\in\mathcal{S}$, action $a\in\mathcal{A}$, and discrete step $n\geq 0$, let $R_{n}(s,a)\in\mathcal{R}$ denote a sample of the resulting reward, and let $S^{\prime}_{n}(s,a)\sim p(\cdot,\cdot\mid s,a)$ denote a sample of the resulting state. Let $\{Y_{n}\}$ be a set-valued process taking values in the set of nonempty subsets of $\mathcal{S}\times\mathcal{A}$, such that: $Y_{n}=\{(s,a):(s,a)$ component of the $|\mathcal{S}\times\mathcal{A}|$-sized table of state-action value estimates, $Q$, that was updated at step $n\}$. Let $\nu(n,s,a)\doteq\sum_{j=0}^{n}I\{(s,a)\in Y_{j}\}$, where $I$ is the indicator function, such that $\nu(n,s,a)$ represents the number of times that the $(s,a)$ component of $Q$ was updated up until step $n$.
Now consider the following update rule for $n\geq 0$:
$Q_{n+1}(s,a)$ $\doteq Q_{n}(s,a)+\alpha_{\nu(n,s,a)}\delta_{n}(s,a)I\{(s,a)\in Y_{n}\},\quad\forall s\in\mathcal{S},a\in\mathcal{A},$ (B.43)
where,
$\begin{split}\delta_{n}(s,a)&\doteq R_{n}(s,a)-\bar{R}_{n}+\max_{a^{\prime}}Q_{n}(S_{n}^{\prime}(s,a),a^{\prime})-Q_{n}(s,a).\end{split}$ (B.44)
Here, $\bar{R}_{n}$ denotes the estimate of the average-reward induced by following policy $\pi$, $\bar{r}_{\pi}$ (see Equation (2)), $\delta_{n}(s,a)$ denotes the TD error, and $\alpha_{\nu(n,s,a)}$ denotes the step size at time step $n$ for state-action pair $(s,a)$.
We will now show that the value function update rule for the D2 Q-learning algorithm is a special case of the update rule (B.43). To this end, consider a sequence of experience from our MDP $\mathcal{M}$: $S_{t},A_{t},R_{t+1},S_{t+1},\ldots$  . Now recall the set-valued process $\{Y_{n}\}$. If we let $n$ = time step $t$, we have:
$Y_{t}(s,a)=\begin{cases}1,s=S_{t}\text{ and }a=A_{t},\\ 0,\text{ otherwise,}\end{cases}$
as well as $S^{\prime}_{n}(S_{t},A_{t})=S_{t+1}$, and $R_{n}(S_{t},A_{t})=R_{t+1}$.
Hence, the update rule (B.43) becomes:
$Q_{t+1}(S_{t},A_{t})$ $\doteq Q_{t}(S_{t},A_{t})+\alpha_{\nu(t,S_{t},A_{t})}\delta_{t}\text{ \; and \; }Q_{t+1}(s,a)\doteq Q_{t}(s,a),\forall s\neq S_{t},a\neq A_{t},$ (B.45)
which is D2 Q-learning’s value function update rule with $\alpha_{\nu(t,S_{t},A_{t})}$ denoting the step size at time $t$.
As such, if we can show that the update rule (B.43) converges, then it would imply that the value function update rule for the D2 Q-learning algorithm also converges. To this end, we will now specify the assumptions that are needed to ensure convergence:
Assumption B.2.10 (Communicating Assumption).
The MDP has a single communicating class. That is, each state in the MDP is accessible from every other state under some deterministic stationary policy.
Assumption B.2.11 (State-Action Value Function Uniqueness).
There exists a unique solution of $q$ only up to a constant in the Bellman equation (5).
Assumption B.2.12 (Asynchronous Step Size Assumption 3).
There exists $\Delta>0$ such that
$\liminf_{n\to\infty}\frac{\nu(n,s,a)}{n+1}\geq\Delta,$
a.s., for all $s\in\mathcal{S},a\in\mathcal{A}$.
Furthermore, for all $x>0$, and
$N(n,x)=\min\Bigg{\{}m>n:\sum_{i=n+1}^{m}\alpha_{i}\geq x\Bigg{\}},$
the limit
$\lim_{n\to\infty}\frac{\sum_{i=\nu(n,s,a)}^{\nu(N(n,x),s,a)}\alpha_{i}}{\sum_{i=\nu(n,s^{\prime},a^{\prime})}^{\nu(N(n,x),s^{\prime},a^{\prime})}\alpha_{i}}$
exists a.s. for all $s,s^{\prime},a,a^{\prime}$.
We are now ready to state the convergence theorem:
Theorem B.2.13 (Convergence of D2 Q-learning Value Function Estimates).
If Assumptions B.2.2, B.2.3, B.2.5, B.2.10, B.2.11, and B.2.12 hold, then the update rule (B.43) converges, almost surely, $Q_{n}\to q_{\infty}\text{ as }n\to\infty$, thereby implying that the D2 Q-learning value function estimates (B.45) converge almost surely to $q_{\infty}\text{ as }n\to\infty\textbf{}$, where $q_{\infty}$ denotes the solution to the Bellman optimality equation (5).
We prove this theorem in the following section. To do so, use a two-timescales argument, such that we leverage Theorem 2 in Section 6 of Borkar [2009], along with the results from Theorems 4.5 and B.1.10.
B.2.4 Proof of Theorem B.2.13
To begin, let us consider Assumption B.2.5. In particular, $\alpha_{n}/(\eta_{{}_{\theta}}\alpha_{n})\to 0$ implies that the quantile step size, $\eta_{{}_{\theta}}\alpha_{n}$, decreases to 0 at a slower rate than the value function step size, $\alpha_{n}$. This implies that the quantile updates move on a faster timescale compared to the value function update. Hence, as argued in Section 6 of Borkar [2009], the (faster) quantile updates (B.3) view the (slower) value function update (B.43) as quasi-static, while the (slower) value function update views the (faster) quantile updates as nearly equilibrated (as we will discuss below, the existence of such an equilibrium point was shown in Theorem B.1.10).
Importantly, given the results of Lemma 4.4, Theorem B.1.10, and Theorem 4.5, the two-timescales argument allows us to view the average-reward estimate, $\bar{R}_{n}$, as nearly equilibrated at $\bar{r}_{\pi}$. In particular, given Lemma 4.4, we have that $\bar{R}_{n}\doteq\frac{1}{m}\sum_{i=1}^{m}\theta_{i,n}$, where $m$ denotes the number of quantiles, and $\theta_{i,n}$ denotes the $i^{\text{th}}$ quantile estimate at step $n$. Now, given Theorem B.1.10, we now that the quantile estimates converge to the desired quantiles of the limiting per-step reward distribution. As such, by Theorem 4.5, we have that $\bar{R}_{n}\to\bar{r}_{\pi}$. Consequently, by the two-timescales argument, the value function update views the average-reward estimate, $\bar{R}_{n}$, as nearly equilibrated at $\bar{r}_{\pi}$ (that is, we can essentially treat $\bar{r}_{\pi}$ as a known constant).
We now proceed to show the convergence of the value function estimates. To this end, consider the asynchronous value function update (B.43). This update is of the same form as Equation 7.1.2 of Borkar [2009]. As such, to show the convergence of the value function estimates, we can apply the result in Section 7.4 of Borkar [2009], which shows the convergence of asynchronous updates of the same form as Equation 7.1.2. To apply this result, given Assumptions B.2.3 and B.2.12, we only need to show the convergence of the synchronous version of the value function update:
$Q_{n+1}=Q_{n}+\alpha_{n}(h(Q_{n})+M_{n+1}),\quad\forall s\in\mathcal{S},a\in\mathcal{A},$ (B.46)
where,
$h(Q_{n})(s,a)$ $\doteq\sum_{s^{\prime},r}p(s^{\prime},r\mid s,a)(r+\max_{a^{\prime}}Q_{n}(s^{\prime},a^{\prime}))-Q_{n}(s,a)-\bar{r}_{\pi},$
$=T(Q_{n})(s,a)-Q_{n}(s,a)-\bar{r}_{\pi},\text{ and}$
$M_{n+1}(s,a)$ $\doteq R_{n}(s,a)+\max_{a^{\prime}}Q_{n}(S_{n}^{\prime}(s,a),a^{\prime})-Q_{n}(s,a)-\bar{r}_{\pi}-h(Q_{n})(s,a).$
To show the convergence of the synchronous update (B.46) under the two-timescales argument, we can apply the result of Theorem 2 in Section 6 of Borkar [2009] to show that $Q_{n}\to q_{\infty}$ a.s. as $n\to\infty$. This theorem requires that 3 assumptions be satisfied. As such, we will now show, via Lemmas B.2.14 - B.2.16, that these 3 assumptions are indeed satisfied.
Lemma B.2.14.
The quantile update rules, which can be written in the following form: $\theta_{i,n+1}=\theta_{i,n}+\eta_{{}_{\theta}}\alpha_{n}(g(\theta_{i,n})+M_{i,n})$ (see Lemma B.1.7), each have a globally asymptotically stable equilibrium, $\theta_{i},\;\forall i=1,2,\ldots m$, where $m$ denotes the number of quantiles.
Proof.
This was shown in Theorem B.1.10. ∎
Lemma B.2.15.
The value function update rule (B.46) has a globally asymptotically stable equilibrium, $q_{\infty}$.
Proof.
The value function update rule (B.46) is of the same form as the update rule considered in Lemmas B.1 and B.2 of Wan et al. [2021]. Given that we adhere to all the assumptions required for the aforementioned Lemmas from Wan et al. [2021], we can leverage their results, which show that the value function update rule (B.46) has a globally asymptotically stable equilibrium, $q_{\infty}$. ∎
Lemma B.2.16.
$\sup_{n}(||Q_{n}||+\sum_{i=1}^{m}{||\theta_{i,n}||)}<\infty$ a.s.
Proof.
It was shown in Lemma B.1.8 that $\sup_{n}(||\theta_{i,n}||)<\infty\;\forall i=1,2,\ldots,m$ a.s. Hence, we only need to show that $\sup_{n}(||Q_{n}||)<\infty$ a.s. To this end, we can apply Theorem 7 in Section 3 of Borkar [2009]. This theorem requires 4 assumptions:
  • (A1) The function $h$ is Lipschitz. That is, $||h(x)-h(y)||\leq L||x-y||$ for some $0<L<\infty$.
  • (A2) The sequence $\{\alpha_{n}\}$ satisfies $\alpha_{n}>0$, $\sum\alpha_{n}=\infty$, and $\sum\alpha_{n}^{2}<\infty$.
  • (A3) $\{M_{n}\}$ is a martingale difference sequence that is square-integrable.
  • (A4) The functions $h_{d}(x)\doteq h(dx)/d$, $d\geq 1,x\in\mathbb{R}^{k}$, satisfy $h_{d}(x)\to h_{\infty}(x)$ as $d\to\infty$, uniformly on compacts for some $h_{\infty}\in C(\mathbb{R}^{k})$. Furthermore, the ODE $\dot{x}_{t}=h_{\infty}(x_{t})$ has the origin as its unique globally asymptotically stable equilibrium.
We first note that Assumption (A1) is satisfied given that both the $\sum$ and $T$ operators are Lipschitz. Moreover, we note that Assumption B.2.2 satisfies Assumption (A2).
We now show that Assumption (A3) is satisfied:
Let $\mathcal{F}_{n}\doteq\sigma(Q_{i},M_{i},i\leq n),n\geq 0$ denote an increasing family of $\sigma$-fields. We have that, for any $s\in\mathcal{S}$:
$\mathbb{E}[M_{n+1}(s,a)\mid\mathcal{F}_{n}]$ $=\mathbb{E}\left[R_{n}(s,a)+\max_{a^{\prime}}Q_{n}(S_{n}^{\prime}(s,a),a^{\prime})-Q_{n}(s,a)-\bar{r}_{\pi}-h(Q_{n})(s,a)\mid\mathcal{F}_{n}\right]$ (B.47)
$=\mathbb{E}\left[R_{n}(s,a)+\max_{a^{\prime}}Q_{n}(S_{n}^{\prime}(s,a),a^{\prime})-Q_{n}(s,a)-\bar{r}_{\pi}\mid\mathcal{F}_{n}\right]-h(Q_{n})(s,a)$ (B.48)
$=\mathbb{E}\left[R_{n}(s,a)+\max_{a^{\prime}}Q_{n}(S_{n}^{\prime}(s,a),a^{\prime})\mid\mathcal{F}_{n}\right]-Q_{n}(s,a)-\bar{r}_{\pi}-h(Q_{n})(s,a)$ (B.49)
$=\mathbb{E}[R_{n}(s,a)+\max_{a^{\prime}}Q_{n}(S_{n}^{\prime}(s,a),a^{\prime})\mid\mathcal{F}_{n}]-T(Q_{n})(s,a)$ (B.50)
$=0.$ (B.51)
Moreover, it can be easily shown via the triangle inequality that $\mathbb{E}[||M_{n+1}||^{2}\mid\mathcal{F}_{n}]\leq K(1+||Q_{n}||^{2})$ for some constant $K>0$. As such, Assumption (A3) is satisfied.
Finally, we show that Assumption (A4) is also satisfied:
We first note that:
$h_{\infty}(x)=\lim_{d\to\infty}h_{d}(x)=\lim_{d\to\infty}\frac{T(dx)-dx-\bar{r}_{\pi}e}{d}=T_{0}(x)-x,$ (B.52)
where
$T_{0}(x)\doteq\sum_{s^{\prime},r}p(s^{\prime},r\mid s,a)\max_{a^{\prime}}x(s^{\prime},a^{\prime}).$ (B.53)
Clearly, the function $h_{\infty}$ is continuous in every $x\in\mathbb{R}^{k}$. As such, we have that $h_{\infty}\in C(\mathbb{R}^{k})$.
Next, consider the ODE $\dot{x}_{t}=h_{\infty}(x_{t})=T_{0}(x_{t})-x_{t}$. This ODE has the origin as an equilibrium since $T_{0}(0)-0=0$. Moreover, we can again leverage the results of Lemmas B.1 and B.2 from Wan et al. [2021] to conclude that the origin is the unique globally asymptotically stable equilibrium for this ODE. Hence, Assumption A4 is satisfied.
As such, Assumptions (A1) - (A4) are all verified, meaning that we can apply the results of Theorem 7 in Section 3 of Borkar [2009] to conclude that $\sup_{n}(||Q_{n}||)<\infty$ a.s., and hence, that $\sup_{n}(||Q_{n}||+\sum_{i=1}^{m}{||\theta_{i,n}||)}<\infty$ a.s. ∎
As such, we have now verified the 3 assumptions required by Theorem 2 in Section 6 of Borkar [2009], which means that we can apply the result of the theorem to conclude that $Q_{n}\to q_{\infty}$ a.s. as $n\to\infty$.
This completes the proof of Theorem B.2.13.
Appendix C Numerical Experiments
This appendix contains details regarding the numerical experiments performed as part of this work. We discuss the experiments performed in the red-pill blue-pill environment Rojas and Lee [2025], and the well-known inverted pendulum environment.
Figure C.1: An illustration of the a) red-pill blue-pill, and b) inverted pendulum environments.
The aim of the experiments was to contrast and compare the Differential Distributional RL algorithms (see Appendix A) with non-distributional average-reward RL algorithms. In particular, we aimed to show how the Differential Distributional RL algorithms could be utilized to find the average-reward-optimal policy (while also learning the corresponding optimal per-step reward distribution), and contrast the results to those of non-distributional algorithms, which served as a sort of ‘baseline’ to illustrate how our distributional approach compares to a non-distributional one. In this work, we utilized non-distributional average-reward RL algorithms derived from the Differential framework proposed in Wan et al. [2021]. The aforementioned (non-distributional) algorithms used for comparison are included in Section C.3.
C.1 Red-Pill Blue-Pill Experiment
The first experiment performed was in the two-state red-pill blue-pill environment, where at every time step, an agent can take either a ‘red pill’, which takes them to the ‘red world’ state, or a ‘blue pill’, which takes them to the ‘blue world’ state. Each state has its own characteristic reward distribution, such that the blue world state has a reward distribution with a higher (better) average-reward compared to the red world state. Hence, we would expect that the D2 and D3 Q-learning algorithms learn a policy that prefers to stay in the blue world state, as well as the limiting per-step reward distribution induced by such a policy. Similarly, we would expect that the Differential Q-learning algorithm learns a policy that prefers to stay in the blue world state. This task is illustrated in Figure C.1a).
For this experiment, we compared our tabular D2 and D3 Q-learning algorithms (Algorithms 4 and 9, respectively) to the tabular Differential Q-learning algorithm (Algorithm 13). We ran each algorithm using various combinations of step sizes for each algorithm. We used an $\varepsilon$-greedy policy with a fixed epsilon of 0.1. We used 10 per-step reward quantiles, and 10 differential return quantiles (i.e., $m=n=10$). We set all initial guesses to zero. We ran the algorithms for 100k time steps.
[…]
Figure C.2: Step size tuning results for the red-pill blue-pill task when using the Differential Q-learning algorithm. The average-reward for all time steps is displayed for various combinations of step size parameters.
[…]
Figure C.3 shows the agent’s (per-step) reward quantile estimates as learning progresses when using the D3 Q-learning algorithm in the red-pill blue-pill task. As shown in the figure, the agent’s quantile estimates converge to the quantiles of the limiting per-step reward distribution induced by the optimal policy (i.e., the policy that remains in the blue world state). A differential return quantile step size of 2e-4 and a per-step reward quantile $\eta$ of 2.0 were used to generate the results displayed in Figure C.3.
Figure C.3: a) Histogram showing the empirical ($\varepsilon$-greedy) optimal (long-run) per-step reward distribution in the red-pill blue-pill task. b) Quantiles of the ($\varepsilon$-greedy) optimal (long-run) per-step reward distribution in the red-pill blue-pill task. c) Convergence plot of the agent’s (per-step) reward quantile estimates as learning progresses when using the D3 Q-learning algorithm in the red-pill blue-pill task.
Figure C.4 shows the agent’s differential return quantile estimates for $S_{t}=$ ‘blue world’ and $A_{t}=$ ‘blue pill’ as learning progresses when using the D3 Q-learning algorithm in the red-pill blue-pill task. As shown in the figure, the agent’s differential return quantile estimates converge. A differential return quantile step size of 2e-4 and a per-step reward quantile $\eta$ of 2.0 were used to generate the results displayed in Figure C.4.
Figure C.4: Convergence plot of the agent’s differential return quantile estimates for $S_{t}=$ ‘blue world’ and $A_{t}=$ ‘blue pill’ as learning progresses when using the D3 Q-learning algorithm in the red-pill blue-pill task.
C.2 Inverted Pendulum Experiment
The second experiment performed was in the well-known inverted pendulum environment, where an agent learns how to optimally balance an inverted pendulum. We chose this task because it provides us with the opportunity to test our algorithms in an environment where: 1) we must use function approximation (given the high-dimensional state-space), and 2) where the optimal policy is known (i.e., the optimal policy is the policy that successfully balances the pendulum, thereby yielding a limiting per-step reward distribution $\approx 0$). This hence allows us to directly assess the quality of the quantile estimates (to see if they all converge to $\approx 0$), as well as to gauge how function approximation affects the performance of our algorithms. For this task, we utilized a simple actor-critic architecture [Barto et al., 1983, Sutton and Barto, 2018] as this allowed us to compare the performance of (non-tabular) D2 and D3 TD-learning algorithms with a (non-tabular) Differential TD-learning algorithm. This task is illustrated in Figure C.1b).
For this experiment, we compared our D2 and D3 Actor-Critic algorithms (Algorithms 7 and 12, respectively) to the Differential Actor-Critic algorithm (Algorithm 14). We ran each algorithm using various combinations of step sizes for each algorithm. We used 10 per-step reward quantiles, and 10 differential return quantiles (i.e., $m=n=10$). We used a quantile Huber loss $\lambda$ of 1.0. We set all initial guesses to zero. We ran the algorithms for 10k time steps. For simplicity, we used tile coding [Sutton and Barto, 2018] for both the value function and policy parameterizations, where we parameterized a softmax policy. For each parameterization, we used 32 tilings, each with 8 X 8 tiles. By using a linear function approximator (i.e., tile coding), the gradients for the value function and policy parameterizations can be simplified as follows:
$\nabla\hat{v}(s,\boldsymbol{w})=\boldsymbol{x}(s),$ (C.1)
$\nabla\text{ln}\pi(a\mid s,\boldsymbol{\theta})=\boldsymbol{x}_{h}(s,a)-\sum_{\xi\in\mathcal{A}}\pi(\xi\mid s,\boldsymbol{\theta})\boldsymbol{x}_{h}(s,\xi),$ (C.2)
where $s\in\mathcal{S}$, $a\in\mathcal{A}$, $\boldsymbol{x}(s)$ is the state feature vector, and $\boldsymbol{x}_{h}(s,a)$ is the softmax preference vector.
[…]
Figure C.5 shows the agent’s (per-step) reward quantile estimates as learning progresses when using the D3 actor-critic algorithm in the inverted pendulum task. As shown in the figure, the agent’s quantile estimates converge to the quantiles of the limiting per-step reward distribution induced by the optimal policy (i.e., the policy that balances the pendulum, thereby yielding a limiting per-step reward distribution $\approx 0$).
Figure C.5: a) Histogram showing the empirical ($\varepsilon$-greedy) optimal (long-run) per-step reward distribution in the inverted pendulum task. b) Quantiles of the ($\varepsilon$-greedy) optimal (long-run) per-step reward distribution in the inverted pendulum task. c) Convergence plot of the agent’s (per-step) reward quantile estimates as learning progresses when using the D3 actor-critic algorithm in the inverted pendulum task.
C.3 Baselines
Below is the pseudocode for the non-distributional baseline algorithms used for comparison in our experiments. We first present the algorithms derived from the Differential framework proposed in Wan et al. [2021].
Algorithm 13 Differential Q-Learning (Tabular)
  Input: the policy $\pi$ to be used (e.g., $\varepsilon$-greedy)
  Algorithm parameters: step size parameters $\alpha$, $\eta$
  Initialize $Q(s,a)\>\forall s,a$ (e.g. to zero)
  Initialize $\bar{R}$ arbitrarily (e.g. to zero)
  Obtain initial $S$
  while still time to train do
     $A\leftarrow$ action given by $\pi$ for $S$
     Take action $A$, observe $R,S^{\prime}$
     $\delta=R-\bar{R}+\max_{a}Q(S^{\prime},a)-Q(S,A)$
     $\bar{R}=\bar{R}+\eta\alpha\delta$
     $Q(S,A)=Q(S,A)+\alpha\delta$
     $S=S^{\prime}$
  return $Q$
Algorithm 14 Differential Actor-Critic
  Input: a differentiable state-value function parameterization $\hat{v}(s,\boldsymbol{w})$, a differentiable policy parameterization $\pi(a\mid s,\boldsymbol{u})$
  Algorithm parameters: step size parameters $\alpha$, $\eta_{\pi}$, $\eta_{{}_{\bar{R}}}$
  Initialize state-value weights $\boldsymbol{w}\in\mathbb{R}^{d}$ and policy weights $\boldsymbol{u}\in\mathbb{R}^{d^{\prime}}$ (e.g. to $\boldsymbol{0}$)
  Initialize $\bar{R}$ arbitrarily (e.g. to zero)
  Obtain initial $S$
  while still time to train do
     $A\sim\pi(\cdot\mid S,\boldsymbol{u})$
     Take action $A$, observe $R,S^{\prime}$
     $\delta=R-\bar{R}+\hat{v}(S^{\prime},\boldsymbol{w})-\hat{v}(S,\boldsymbol{w})$
     $\bar{R}=\bar{R}+\eta_{{}_{\bar{R}}}\alpha\delta$
     $\boldsymbol{w}=\boldsymbol{w}+\alpha\delta\nabla\hat{v}(S,\boldsymbol{w})$
     $\boldsymbol{u}=\boldsymbol{u}+\eta_{\pi}\alpha\delta\nabla\text{ln}\pi(A\mid S,\boldsymbol{u})$
     $S=S^{\prime}$
  return $\boldsymbol{w}$ and $\boldsymbol{u}$